### v1

# Continuous-Aperture Array (CAPA)-Based Wireless Communications: Capacity Characterization

Abstract
The capacity limits of continuous-aperture array (CAPA)-based wireless communications are characterized. To this end, an analytically tractable transmission framework is established for both uplink and downlink CAPA systems. Based on this framework, closed-form expressions for the single-user channel capacity are derived. The results are further extended to a multiuser case by characterizing the capacity limits of a two-user channel and proposing the associated capacity-achieving decoding and encoding schemes. 1) For the uplink case, the sum-rate capacity and capacity region, as well as the capacity-achieving detectors, are derived. 2) For the downlink case, the uplink-downlink duality is established by deriving the uplink-to-downlink and downlink-to-uplink transformations under the same power constraint, based on which the optimal power allocation policy and the achieved sum-rate capacity and capacity region are characterized. To gain further insights, several case studies are presented by specializing the derived results into various array structures, including the planar CAPA, linear CAPA, and planar spatially discrete array (SPDA). Numerical results are provided to reveal that: i) the channel capacity achieved by CAPAs converges towards a finite upper bound as the aperture size increases; and ii) CAPAs offer significant capacity gains over the conventional SPDAs.
Index Terms:
Channel capacity, continuous aperture array (CAPA), performance analysis.
I Introduction
Multiple antenna technology serves as a cornerstone in the advancement of modern wireless communication systems. At its core lies the principle of leveraging an increased number of antenna elements to enhance spatial degrees of freedom (DoFs) and augment channel capacity [2]. Traditionally, multiple antenna prototypes follow a spatially discrete topological structure, where each antenna is depicted as a distinct point [3]. Motivated by the advantages of incorporating a greater number of antennas, the concept of densely packed antennas has garnered considerable interest in the communication field.
Within a fixed array aperture, reducing the distance between elements enables the addition of more antennas, thus increasing spatial DoFs. This development has led to the emergence of holographic arrays [4, 5]. In such arrays, antenna elements are densely packed in a configuration that is spaced less than half a wavelength apart, resulting in enhanced spectral efficiency [6].
Building on the advancements discussed, this paper introduces a novel concept in antenna design—continuous aperture array (CAPA). CAPA represents a paradigm shift from the conventional spatially discrete array (SPDA), evolving towards an array that forms a spatially continuous electromagnetic (EM) aperture. Essentially, a CAPA can be envisioned as an SPDA comprising an infinite number of antennas with infinitesimal spacing [3, 7]. The concept of a CAPA marks a significant advancement towards maximizing the spatial DoFs available in antenna array configurations, thereby providing substantial gains in system performance. Transitioning from discrete to continuous aperture arrays enhances signal resolution and provides more flexible control over EM wave propagation. Through this innovative approach, CAPA paves the way for next-generation antenna technologies, promising to revolutionize wireless communication infrastructures with its advanced spatial processing capabilities [8].
However, unlike traditional arrays that are modeled using spatially discrete approaches, the modeling of CAPA-based wireless transmission introduces a fundamentally different method grounded in EM field theory. Specifically, while the channels of conventional SPDAs are typically represented through discrete matrices, the spatial response of a CAPA necessitates representation through a continuous operator in Hilbert space [9, 10]. This shift from discrete to continuous modeling is not merely a technical detail but a transformative change in how wireless transmission systems are understood and designed [11]. The use of continuous operators to model the spatial responses of CAPAs signifies a departure from the traditional matrix-based methods suited to SPDAs. This change is crucial because it implies that the entire framework for analyzing and designing transmissions in traditional SPDAs becomes fundamentally inapplicable to CAPAs. Consequently, new conceptual and mathematical frameworks tailored specifically to handle the continuous nature of EM field interactions within CAPAs are necessary [9, 10]. This development is essential for fully exploiting the potential of CAPAs in enhancing system performance.
Currently, research on CAPA remains in a nascent stage and is predominantly focused on single-user scenarios [11, 12, 13, 14, 15]. While these foundational studies have laid the groundwork, exploration into the performance of multiuser CAPA systems is considerably less developed. Their performance, such as channel capacity, has been seldom analyzed, with only [16] presenting some initial results. Specifically, the authors of [16] proposed a sophisticated channel model that utilizes a Green’s function method and leverages a continuous electric current distribution. Their research also extends to the performance analysis of CAPA-based reconfigurable intelligent surfaces (RIS) and simultaneously transmitting and reflecting RISs, which provides new insights into the capabilities of CAPAs in more dynamic settings.
Despite these advancements, a significant gap remains in our understanding of the full potential of CAPAs, particularly regarding their fundamental performance limits. More importantly, an analytically tractable framework for CAPA-based communications is lacking. This framework is essential for advancing both the theoretical understanding and practical deployment of these systems. It would facilitate a deeper comprehension of CAPAs’ capabilities and enable the design of more effective and efficient communication strategies tailored to their unique properties, especially in multiuser scenarios where complexity and demand for system efficiency are significantly higher.
Motivated by these knowledge gaps, this paper proposes a novel analytical framework to address the continuous nature of EM fields within CAPAs, providing a robust foundation for modeling and understanding CAPA-based communications. Using this framework, the channel capacity for both downlink and uplink CAPA-based communications is characterized. The main contributions are summarized as follows.
  • We propose a transmission framework for uplink and downlink CAPA communications using EM field theories, where continuous operators model both signals and spatial responses. Based on this framework, we design optimal detectors and source current distributions to achieve the single-user channel capacity for uplink and downlink CAPA communications, and derive closed-form expressions for the corresponding channel capacity.
  • We then extend our analysis to multiuser CAPA systems by studying a two-user case. For uplink CAPA communications, we propose using successive decoding to achieve capacity and derive closed-form expressions for the optimal detectors of each user. Based on this, we derive closed-form expressions for the per-user rate and the sum-rate capacity, and characterize the capacity region. For downlink CAPA communications, we establish the uplink-downlink duality by presenting the uplink-to-downlink and downlink-to-uplink transformations under the same sum power constraint and derive closed-form expressions for the capacity-achieving source current distributions. Additionally, we characterize the downlink capacity limits by deriving the per-user rate, sum-rate capacity, and capacity region.
  • To unveil further system insights, we present several case studies exploring the line-of-sight (LoS) channel model. The derived results are specialized to various array structures, including the planar CAPA, linear CAPA, and planar SPDA. We derive closed-form expressions for the sum-rate capacity in each case and perform asymptotic analyses by setting the aperture size to infinity.
  • We present numerical results illustrating that the capacity achieved by the CAPA increases with aperture size and converges to a finite upper limit. Our results demonstrate that, compared to CAPA, the capacity of the conventional SPDA is limited by its array occupation ratio. Moreover, CAPA offers significant capacity enhancements over conventional SPDA, whose capacity converges to that of CAPA as the array occupation ratio approaches one.
The remainder of this paper is organized as follows. Section II introduces the transmission framework for CAPA-based communications. Sections III and IV delve into the capacity analysis and the capacity-achieving schemes for single-user and multiuser CAPA systems, respectively. Section V provides case studies for several specific array structures. Section VI presents numerical results to demonstrate the superior performance of CAPAs. Finally, Section VII concludes the paper.
Figure 1: Illustration of CAPA-based communications.
II System Model
[…]
II-A Signal Model
II-A1 Uplink Communications
[…]
${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})$ $=\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){x}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}}$ (1a)
$=s_{{\mathsf{ul}},k}\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){\mathsf{J}}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}},$ (1b)
[…]
${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})\approx\mathsf{H}_{k}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},k}s_{{\mathsf{ul}},k}\int_{{\mathcal{A}}_{k}}{\rm{d}}{\mathbf{s}}=\mathsf{H}_{k}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},k}A_{{\mathsf{u}},k}s_{{\mathsf{ul}},k}.$ (2)
[…]
${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ $=\sum\nolimits_{k=1}^{K}{\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ (3a)
$\approx\sum\nolimits_{k=1}^{K}\mathsf{H}_{k}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},k}A_{{\mathsf{u}},k}s_{{\mathsf{ul}},k}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}),$ (3b)
where ${\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ accounts for thermal noise [18]. The noise field is modeled as a zero-mean complex Gaussian random process satisfying ${\mathbbmss{E}}\{{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}^{*}({\mathbf{r}}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-{\mathbf{r}}^{\prime})$ [18], where $\delta(\cdot)$ represents the Dirac delta function, and $\sigma^{2}$ describes the noise intensity. ${\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ and $\{s_{{\mathsf{ul}},k}\}_{k=1}^{K}$ are assumed to be uncorrelated. After observing ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$, the BS should employ a properly designed detector along with maximum-likelihood (ML) decoding to recover the data information encoded in $\{s_{{\mathsf{ul}},k}\}_{k=1}^{K}$, which will be detailed in Section III-A.
II-A2 Downlink Communications
At the BS, the downlink electrical signal is designed as follows:
${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum\nolimits_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k}$ (4)
[…]
${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})=\int_{{\mathcal{A}}}{\mathsf{H}}(\mathbf{s},{\mathbf{r}})\left(\sum\nolimits_{k^{\prime}=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k^{\prime}}({\mathbf{r}})s_{{\mathsf{dl}},k^{\prime}}\right){\rm{d}}{\mathbf{r}}.$ (5)
Since the aperture size of each user is generally negligible with respect to both the propagation distance and the BS aperture size, the variations of the electric field and channel response along the receive aperture can be omitted. This yields ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s}_{k})\triangleq{\mathsf{E}}_{{\mathsf{dl}},k}$ and ${\mathsf{H}}(\mathbf{s},{\mathbf{r}})\approx{\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})={\mathsf{H}}({\mathbf{r}},\mathbf{s}_{k})=\mathsf{H}_{k}(\mathbf{r})$ for $\mathbf{s}\in{\mathcal{A}}_{k}$. Therefore, we obtain
${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}=\!\int_{{\mathcal{A}}}{\mathsf{H}}_{k}({\mathbf{r}})\left(\sum\nolimits_{k^{\prime}=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k^{\prime}}({\mathbf{r}})s_{{\mathsf{dl}},k^{\prime}}\!\right){\rm{d}}{\mathbf{r}}.$ (6)
Accordingly, the observation of user $k$ at point $\mathbf{s}\in{\mathcal{A}}_{k}$ can be written as follows:
${\mathsf{Y}}_{{\mathsf{dl}},k}(\mathbf{s})$ $={\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})+{\mathsf{N}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}+{\mathsf{N}}_{{\mathsf{dl}},k}(\mathbf{s})$ (7a)
$=s_{{\mathsf{dl}},k}\int_{{\mathcal{A}}}{\mathsf{H}}_{k}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}}){\rm{d}}{\mathbf{r}}+{\mathsf{N}}_{{\mathsf{dl}},k}(\mathbf{s})$
$+\sum\nolimits_{k^{\prime}\neq k}s_{{\mathsf{dl}},k^{\prime}}\int_{{\mathcal{A}}}{\mathsf{H}}_{k}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},k^{\prime}}({\mathbf{r}}){\rm{d}}{\mathbf{r}},$ (7b)
where ${\mathsf{N}}_{{\mathsf{dl}},k}(\mathbf{s})$ denotes the thermal noise subject to ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}},k}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}},k}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{k}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$.
II-B Channel Model
To facilitate theoretical investigations into fundamental performance limits and asymptotic behaviors, we focus our discussion on LoS channels. Under this consideration, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ is modeled as follows [19]:
${\mathsf{H}}(\mathbf{r},{\mathbf{s}})={\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}}){\mathsf{H}}^{\mathsf{pa}}(\mathbf{r},{\mathbf{s}}),$ (8)
[…]
${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\triangleq\frac{{\rm{j}}k_{0}\eta{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}\left(1+\frac{{\rm{j}}k_{0}^{-1}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}\right)$ (9)
models the influence of free-space EM propagation, where $\eta=120\pi$ is the impedance of free space, and $k_{0}=\frac{2\pi}{\lambda}$ with $\lambda$ being the wavelength denotes the wavenumber. Particularly, ${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})$ comprises three terms: the first term corresponds to the radiating near-field and far-field regions, while the remaining two terms correspond to the reactive near-field region. The squared magnitude of the parenthesis in (9) is $1-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}+\frac{k_{0}^{-4}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{4}}$, which equals $0.97$ at distance $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert=\lambda$ [19]. Hence, when considering practical systems with $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert\gg\lambda$, the last two terms in (9) can be neglected. As a result, (9) is approximated as follows:
${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\approx\frac{{\rm{j}}k_{0}\eta}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}.$ (10)
III Single-User Case
[…]
III-A Uplink Capacity
We commence with the uplink scenario. In the single-user case, (3b) simplifies to
${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})=\mathsf{H}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}}}A_{\mathsf{u}}s_{{\mathsf{ul}}}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}).$ (11)
We then design a detector ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to recover $s_{{\mathsf{ul}}}$ from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$. Specifically, applying this detector to ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ yields
$\begin{split}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}&=s_{{\mathsf{ul}}}{\mathsf{J}}_{{\mathsf{ul}}}A_{\mathsf{u}}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\\ &+\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}.\end{split}$ (12)
Since ${\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ is a zero-mean complex Gaussian random field, $\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}$ is complex Gaussian distributed. Its mean and variance are calculated as follows.
Lemma 1.
Under the assumption that ${\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ is a zero-mean complex Gaussian process with ${\mathbbmss{E}}\{{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}^{*}({\mathbf{r}}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-{\mathbf{r}}^{\prime})$, we have $\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\sim{\mathcal{CN}}(0,\sigma^{2}\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}})$.
Proof:
Please refer to Appendix -A for more details. ∎
The signal-to-noise ratio (SNR) for decoding $s_{{\mathsf{ul}}}$ is thus given by
$\gamma_{\mathsf{ul}}=\frac{A_{\mathsf{u}}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}}}\rvert^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma^{2}\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}.$ (13)
The subsequent task is to design ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to maximize $\gamma_{\mathsf{ul}}$. As observed from (13), $\gamma_{\mathsf{ul}}$ is independent of the norm of ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$, i.e., $\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}$. Thus, the optimal detector satisfies
${\mathsf{V}}_{\mathsf{ul}}^{\star}(\mathbf{r})$ $\propto\operatorname*{argmax}_{\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}=1}{\left\lvert\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\right\rvert^{2}}=\frac{{\mathsf{H}}({\mathbf{r}})}{{\mathsf{h}}_{\mathsf{u}}^{{1}/{2}}},$ (14)
where ${\mathsf{h}}_{\mathsf{u}}\triangleq\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}$.
Remark 1.
The results in (14) suggest that the capacity-achieving detector for uplink single-user CAPA communications aligns with the spatial channel response ${\mathsf{H}}(\mathbf{r})$, which can be regarded as the continuous version of maximal-ratio combining (MRC).
Substituting (14) back into (13) gives
$\gamma_{\mathsf{ul}}=\frac{A_{\mathsf{u}}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}}}\rvert^{2}}{\sigma^{2}}\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}=\overline{\gamma}_{\mathsf{ul}}\int_{{\mathcal{A}}}\lvert{\mathsf{G}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}},$ (15)
where $\overline{\gamma}_{\mathsf{ul}}=\frac{A_{\mathsf{u}}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}}}\rvert^{2}k_{0}^{2}\eta^{2}}{4\pi\sigma^{2}}$ and ${\mathsf{G}}(\mathbf{r})=\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}}{\sqrt{4\pi}\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}\sqrt{\frac{\lvert{\mathbf{e}}^{\mathsf{T}}({\mathbf{s}}_{\mathsf{u}}-{\mathbf{r}})\rvert}{\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}}$. Note that $\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}}{\sqrt{4\pi}\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}$ and $\sqrt{\frac{\lvert{\mathbf{e}}^{\mathsf{T}}({\mathbf{s}}_{\mathsf{u}}-{\mathbf{r}})\rvert}{\lVert{\mathbf{r}}-{\mathbf{s}}_{\mathsf{u}}\rVert}}$ characterize the impact of spherical-wave propagation and projected aperture in arbitrary homogeneous mediums, respectively. Thus, we define $\int_{{\mathcal{A}}}\lvert{\mathsf{G}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}\triangleq{\mathsf{g}}_{\mathsf{u}}$ as the channel gain. The remaining term $\overline{\gamma}_{\mathsf{ul}}$ can be treated as the uplink transmit SNR. As a result, building upon (15), the single-user uplink channel capacity can be written as follows:
${\mathsf{C}}_{\mathsf{ul}}=\log_{2}(1+\gamma_{\mathsf{ul}})=\log_{2}(1+{\overline{\gamma}_{\mathsf{ul}}}{\mathsf{g}}_{\mathsf{u}}).$ (16)
III-B Downlink Capacity
Turning to the downlink case, for $K=1$, (7b) simplifies to
${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}+{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}),$ (17)
where ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{\mathsf{u}}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$. Applying the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ to ${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})$ yields
$\begin{split}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}&=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\\ &+\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}.\end{split}$ (18)
Similar to Lemma 1, we have $\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}\sim{\mathcal{CN}}(0,{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}})$. As a result, the SNR for decoding $s_{{\mathsf{dl}}}$ is given by
$\gamma_{\mathsf{dl}}=\frac{\lvert\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\rvert^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}}.$ (19)
By noting that ${A_{\mathsf{u}}}$ is negligible with respect to the propagation distance, we have
$\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\approx{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})A_{\mathsf{u}},\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}\approx\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}},$
which simplifies (19) as follows:
$\gamma_{\mathsf{dl}}$ $\approx\frac{\lvert{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}}$ (20a)
$=\frac{A_{\mathsf{u}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}}.$ (20b)
The results in (20b) suggest that the downlink SNR is independent of the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ under our settings. Thus, any detector can be used to recover $s_{{\mathsf{dl}}}$. Additionally, from (20b), it is observed that the SNR is maximized when the current distribution satisfies ${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\propto{\mathsf{H}}^{*}({\mathbf{r}})$, i.e.,
${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})=\frac{{\mathsf{H}}^{*}({\mathbf{r}})}{{\mathsf{h}}_{\mathsf{u}}^{1/2}}\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}},$ (21)
where $\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}$ is proportional to the transmit power.
Remark 2.
The results in (21) suggest that the capacity-achieving source current for downlink single-user CAPA communications aligns with ${\mathsf{H}}({\mathbf{r}})$, which can be regarded as the continuous version of maximal-ratio transmission (MRT).
Inserting (21) into (20b) gives
$\gamma_{\mathsf{dl}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}{{\sigma}_{\mathsf{u}}^{2}}\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}={\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}},$ (22)
where ${\overline{\gamma}_{\mathsf{dl}}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}k_{0}^{2}\eta^{2}}{4\pi{\sigma}_{\mathsf{u}}^{2}}$ can be interpreted as the downlink transmit SNR. Based on (22), the single-user downlink capacity is given by
${\mathsf{C}}_{\mathsf{dl}}=\log_{2}(1+\gamma_{\mathsf{dl}})=\log_{2}(1+{\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}}).$ (23)
IV Multiuser Case
This section extends the single-user scenario to the multiuser one. We characterize the channel capacity for both uplink and downlink communications. For brevity, our analysis focuses on the two-user case, while extensions to scenarios with an arbitrary number of users will be addressed in future work.
IV-A Uplink Capacity
We first analyze the uplink scenario.
IV-A1 Optimal Uplink Detection
When $K=2$, the observed electric field in (3b) is given by
${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})=\mathsf{H}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},1}A_{{\mathsf{u}},1}s_{{\mathsf{ul}},1}+\mathsf{H}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},2}A_{{\mathsf{u}},2}s_{{\mathsf{ul}},2}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}).$ (24)
The capacity of an uplink multiuser channel can be achieved using successive interference cancellation (SIC) decoding [2]. Specifically, the message sent by one user is first decoded by treating the message from the other user as interference. Once decoded, this interference is removed, and the other message is then decoded without inter-user interference (IUI).
Given the uplink signal model in (24), there are two possible SIC orders: $1\rightarrow 2$ and $2\rightarrow 1$. Here, we discuss the channel capacity achieved by the decoding order $2\rightarrow 1$. In this case, $s_{{\mathsf{ul}},2}$ is decoded from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ by treating $\mathsf{H}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},1}A_{{\mathsf{u}},1}s_{{\mathsf{ul}},1}$ as interference. After $s_{{\mathsf{ul}},2}$ is decoded and subtracted from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$, $s_{{\mathsf{ul}},1}$ is decoded without IUI from the following signal:
${\mathsf{Z}}(\mathbf{r})\triangleq\mathsf{H}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},1}A_{{\mathsf{u}},1}s_{{\mathsf{ul}},1}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}),$ (25)
which is the same as that given in (11). Following the derivation steps to obtain (14), the BS can use the detector ${\mathsf{V}}_{{\mathsf{ul}},1}(\mathbf{r})=\frac{{\mathsf{H}_{1}}({\mathbf{r}})}{\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{H}}_{1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}}$ to recover $s_{{\mathsf{ul}},1}$ from (25). Therefore, the SNR for decoding the message sent by user $1$ is given by
$\gamma_{\mathsf{ul},1}=\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1},$ (26)
where $\overline{\gamma}_{\mathsf{ul},k}=\frac{A_{\mathsf{u},k}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}},k}\rvert^{2}k_{0}^{2}\eta^{2}}{4\pi\sigma^{2}}$ denotes the transmit SNR of user $k\in\{1,2\}$, and ${\mathsf{g}}_{k}=\int_{{\mathcal{A}}}\lvert{\mathsf{G}}_{k}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}$ denotes the associated channel gain with ${\mathsf{G}}_{k}(\mathbf{r})=\frac{{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}_{k}\rVert}}{\sqrt{4\pi}\lVert{\mathbf{r}}-{\mathbf{s}}_{k}\rVert}\sqrt{\frac{\lvert{\mathbf{e}}^{\mathsf{T}}({\mathbf{s}}_{k}-{\mathbf{r}})\rvert}{\lVert{\mathbf{r}}-{\mathbf{s}}_{k}\rVert}}$.
Next, we investigate the decoding of $s_{{\mathsf{ul}},2}$ from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$, where the interference-plus-noise term is given by ${\mathsf{Z}}(\mathbf{r})$ as shown in (25). Since ${\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})$ and $s_{{\mathsf{ul}},1}$ are uncorrelated, the autocorrelation function of the random field ${\mathsf{Z}}(\mathbf{r})$ is given by
$\begin{split}{\mathbbmss{E}}\{{\mathsf{Z}}(\mathbf{r}){\mathsf{Z}}^{*}(\mathbf{r}^{\prime})\}&={\mathsf{G}}_{1}(\mathbf{r}){\mathsf{G}}_{1}^{*}(\mathbf{r}^{\prime})\overline{\gamma}_{\mathsf{ul},1}{\sigma}^{2}\\ &+{\sigma}^{2}\delta(\mathbf{r}-{\mathbf{r}}^{\prime})\triangleq{\mathsf{R}}_{{\mathsf{Z}}}(\mathbf{r},{\mathbf{r}}^{\prime}).\end{split}$ (27)
To obtain the optimal detector that maximizes the achievable rate of user $2$, we first need to design an invertible linear transformation ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to whiten ${\mathsf{Z}}(\mathbf{r})$. This transformation ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ must satisfy the following two conditions:
  1. i
    Whitening: The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Z}}(\mathbf{r}){\rm{d}}\mathbf{r}$ is proportional to the Dirac delta function, i.e.,
    ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}\propto\delta(\mathbf{r}^{\prime}-\mathbf{r});$ (28)
  2. ii
    Invertibility: Given an arbitrary function $f(\mathbf{r})$ defined on $\mathbf{r}\in{\mathcal{A}}$, there exists a transformation $\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime})$ such that
    $\int_{{\mathcal{A}}}\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime})\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})f(\mathbf{r}){\rm{d}}\mathbf{r}{\rm{d}}\mathbf{r}^{\prime}=f(\mathbf{x}).$ (29)
Applying an invertible transformation to ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ is information lossless and has no impact on the channel capacity [2]. To proceed, we introduce the following lemmas.
Lemma 2.
The invertible linear transformation used to whiten the interference-plus-noise term ${\mathsf{Z}}(\mathbf{r})$ is given by
${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})=\delta(\mathbf{r}^{\prime}-\mathbf{r})+\mu_{1}{\mathsf{G}}_{1}(\mathbf{r}^{\prime}){\mathsf{G}}_{1}^{*}(\mathbf{r}),$ (30)
where $\mu_{1}=-\frac{1}{{\mathsf{g}}_{1}}\pm\frac{1}{{\mathsf{g}}_{1}\sqrt{1+\overline{\gamma}_{\mathsf{u},1}{\mathsf{g}}_{1}}}$. The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\cdot)$ satisfies ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$. Additionally, the inversion of ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ is given by $\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{r}^{\prime})=\delta(\mathbf{r}-\mathbf{r}^{\prime})-\frac{\mu_{1}}{1+\mu_{1}{\mathsf{g}}_{1}}{\mathsf{G}}_{1}(\mathbf{r}){\mathsf{G}}_{1}^{*}(\mathbf{r}^{\prime})$.
Proof:
Please refer to Appendix -B for more details. ∎
Next, we exploit ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to transform ${\mathsf{Y}}_{\mathsf{B}}(\mathbf{r})$ as follows:
$\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}\mathbf{r}=\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime}){\mathsf{J}}_{{\mathsf{ul}},2}A_{{\mathsf{u}},2}s_{{\mathsf{ul}},2}+{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime}),$ (31)
where $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}\mathbf{r}$. By noting that ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$, the model presented in (31) is similar to that in (11). Motivated by this, we design an MRC-based detector $\mathsf{V}_{\mathsf{ul},2}(\mathbf{r}^{\prime})$ by treating $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ as the equivalent channel response, which yields
${\mathsf{V}}_{\mathsf{ul},2}(\mathbf{r}^{\prime})=\frac{{\overline{\mathsf{H}}_{2}}({\mathbf{r}^{\prime}})}{\sqrt{\int_{{\mathcal{A}}}\lvert\overline{\mathsf{H}}_{2}({\mathbf{r}^{\prime}})\rvert^{2}{\rm{d}}{\mathbf{r}}}}.$ (32)
Following the derivation steps to obtain (15), the SNR for decoding $s_{{\mathsf{ul}},2}$ can be written as follows:
$\begin{split}\gamma_{\mathsf{ul},2}=\frac{A_{{\mathsf{u}},2}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}},2}\rvert^{2}}{\sigma^{2}}\int_{{\mathcal{A}}}\lvert\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})\rvert^{2}{\rm{d}}{\mathbf{r}}^{\prime}.\end{split}$ (33)
A closed-form expression for $\gamma_{\mathsf{ul},2}$ is given as follows.
Theorem 1.
By first whitening ${\mathsf{Z}}(\mathbf{r})$ with ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ and then employing the MRC detector ${\mathsf{V}}_{\mathsf{ul},2}(\mathbf{r}^{\prime})$, the resultant SNR in decoding $x_{\mathsf{u},2}$ can be written as follows:
$\begin{split}\gamma_{\mathsf{ul},2}=\overline{\gamma}_{\mathsf{ul},2}\mathsf{g}_{2}(1-{\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}\lvert\rho\rvert^{2}}({1+\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}})^{-1}),\end{split}$ (34)
where $\rho\triangleq\frac{\int_{{\mathcal{A}}}{\mathsf{G}}_{1}^{*}({\mathbf{r}}){\mathsf{G}}_{2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}}{\sqrt{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}}$ is defined as the channel correlation factor with $\lvert\rho\rvert\in[0,1]$.
Proof:
Please refer to Appendix -C for more details. ∎
After obtaining the capacity-achieving detectors under the SIC order $2\rightarrow 1$, we summarize the entire decoding procedure in Table I for ease of reference. The presented decoding procedure can be directly extended to the other SIC order $1\rightarrow 2$ by exchanging the user indices.
1:Whiten the interference-plus-noise term ${\mathsf{Z}}(\mathbf{r})$: $\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}\mathbf{r}=\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime}){\mathsf{J}}_{{\mathsf{ul}},2}A_{{\mathsf{u}},2}s_{{\mathsf{ul}},2}+{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime});$ 2:Use the MRC detector and ML decoder to recover $s_{{\mathsf{ul}},2}$: $\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul},2}^{*}(\mathbf{r}^{\prime})\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}\mathbf{r}{\rm{d}}\mathbf{r}^{\prime}\rightarrow s_{{\mathsf{ul}},2};$ 3:Employ SIC to subtract $\mathsf{H}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},2}A_{{\mathsf{u}},2}s_{{\mathsf{ul}},2}$ from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$; 4:Use the MRC detector and ML decoder to recover $s_{{\mathsf{ul}},1}$: $\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul},1}^{*}(\mathbf{r})(\mathsf{H}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},1}A_{{\mathsf{u}},1}s_{{\mathsf{ul}},1}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})){\rm{d}}\mathbf{r}\rightarrow s_{{\mathsf{ul}},1}.$
TABLE I: SIC decoding for uplink CAPA communications.
IV-A2 Capacity Characterization
After obtaining the decoding SNR of each $s_{{\mathsf{ul}},k}$ for $k=1,2$, the uplink achievable rate of user $k$ can be calculated as ${\mathsf{R}}_{\mathsf{ul},k}^{2\rightarrow 1}=\log_{2}(1+\gamma_{\mathsf{ul},k})$, i.e.,
${\mathsf{R}}_{\mathsf{ul},1}^{2\rightarrow 1}$ $=\log_{2}(1+\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}),$ (35)
${\mathsf{R}}_{\mathsf{ul},2}^{2\rightarrow 1}$ $=\log_{2}\bigg{(}1+\overline{\gamma}_{\mathsf{ul},2}\mathsf{g}_{2}\bigg{(}1-\frac{\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}\lvert\rho\rvert^{2}}{{1+\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}}}\bigg{)}\bigg{)}.$ (36)
Accordingly, the sum-rate capacity is calculated as follows:
$\begin{split}{\mathsf{C}}_{\mathsf{ul}}^{2\rightarrow 1}&={\mathsf{R}}_{\mathsf{ul},1}^{2\rightarrow 1}+{\mathsf{R}}_{\mathsf{ul},2}^{2\rightarrow 1}\\ &=\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{ul},1}\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}+\sum\nolimits_{k=1}^{2}\overline{\gamma}_{\mathsf{ul},k}{\mathsf{g}}_{k}\Big{)},\end{split}$ (37)
where $\overline{\rho}=1-\lvert{\rho}\rvert^{2}\in[0,1]$. Following the same derivation steps used to obtain (37), we can calculate the sum-rate capacity achieved under the SIC order $1\rightarrow 2$ as follows.
Corollary 1.
Under the SIC order $1\rightarrow 2$, the achievable rates of user 1 and user 2, as well as the sum-rate capacity, can be expressed as follows:
${\mathsf{R}}_{\mathsf{ul},1}^{1\rightarrow 2}=\log_{2}\bigg{(}1+\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}\bigg{(}1-\frac{\overline{\gamma}_{\mathsf{ul},2}\mathsf{g}_{2}\lvert\rho\rvert^{2}}{1+\overline{\gamma}_{\mathsf{ul},2}\mathsf{g}_{2}}\bigg{)}\bigg{)},$ (38)
${\mathsf{R}}_{\mathsf{ul},2}^{1\rightarrow 2}=\log_{2}(1+\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{2}),$ (39)
${\mathsf{C}}_{\mathsf{ul}}^{1\rightarrow 2}=\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{ul},1}\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}+\sum\nolimits_{k=1}^{2}\overline{\gamma}_{\mathsf{ul},k}{\mathsf{g}}_{k}\Big{)}.$ (40)
Comparing (37) with (40) yields the following conclusion.
Theorem 2.
The sum-rate capacity of the considered uplink CAPA-based channel is always the same, i.e.,
${\mathsf{C}}_{\mathsf{ul}}=\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{ul},1}\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}+\sum\nolimits_{k=1}^{2}\overline{\gamma}_{\mathsf{ul},k}{\mathsf{g}}_{k}\Big{)},$ (41)
regardless of the decoding order.
Remark 3.
The sum-rate capacity expression given in (41) is determined by the channel gain of each user and the channel correlation factor. This expression is applicable to an arbitrary aperture regardless of its location, shape, and size. Furthermore, although only the LoS channel model is considered, the above derivations can be directly extended to other channel types, as they are not channel-specific.
Having obtained the sum-rate capacity, we now study the capacity region. The capacity region comprises all achievable rate pairs $({\mathsf{R}}_{1},{\mathsf{R}}_{2})$, where ${\mathsf{R}}_{k}$ denotes the achievable rate of user $k\in\{1,2\}$. The capacity region is defined by the rate pairs such that [2]
${\mathsf{R}}_{1}\leq\log_{2}(1+\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}),{\mathsf{R}}_{2}\leq\log_{2}(1+\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{2}),$ (42)
${\mathsf{R}}_{1}+{\mathsf{R}}_{2}\leq\log_{2}\Big{(}1\!+\!\overline{\gamma}_{\mathsf{ul},1}\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}\!+\!\sum\nolimits_{k=1}^{2}\!\overline{\gamma}_{\mathsf{ul},k}{\mathsf{g}}_{k}\Big{)}.$ (43)
This region can be achieved through SIC decoding along with time sharing, which forms a pentagon.
IV-B Downlink Capacity
In this part, we focus on the downlink scenario.
IV-B1 Optimal Multiuser Precoding
The capacity of a downlink multiuser Gaussian channel can be achieved using dirty-paper coding (DPC) [20]. We consider the encoding order $\varepsilon_{K}\rightarrow\varepsilon_{K-1}\rightarrow\ldots\rightarrow\varepsilon_{1}$ with $\{\varepsilon_{k}\}_{k=1}^{K}=\{1,\ldots,K\}$. For user $\varepsilon_{k}$, the dirty-paper encoder treats the interference from user $\varepsilon_{k^{\prime}}$ for $k^{\prime}>k$ as non-causally known, while its decoder regards the interference from user $\varepsilon_{k^{\prime}}$ for $k^{\prime}<k$ as additional noise. Let ${\mathsf{V}}_{{\mathsf{dl}},k}(\mathbf{s})$ denote the detector used by user $k$ to recover $s_{{\mathsf{dl}},k}$ from (7b). By applying DPC and using ML decoding at each user, the signal-to-noise-plus-interference ratio (SINR) for user $\varepsilon_{k}$ is given by
$\gamma_{\mathsf{dl},\varepsilon_{k}}=\frac{\overline{\mathsf{v}}_{\varepsilon_{k}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{\varepsilon_{k}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{\varepsilon_{k}}^{2}{\mathsf{v}}_{\varepsilon_{k}}+\sum_{k^{\prime}<k}\overline{\mathsf{v}}_{\varepsilon_{k}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{\varepsilon_{k}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},\varepsilon_{k^{\prime}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}},$ (44)
where ${\mathsf{v}}_{\varepsilon_{k}}=\int_{\mathcal{A}_{\varepsilon_{k}}}{\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}(\mathbf{s})\rvert^{2}}\mathrm{d}\mathbf{s}$ and $\overline{\mathsf{v}}_{\varepsilon_{k}}=\lvert\int_{\mathcal{A}_{\varepsilon_{k}}}{\mathsf{V}}_{{\mathsf{dl}},\varepsilon_{k}}^{*}(\mathbf{s})\mathrm{d}\mathbf{s}\rvert^{2}$. Since each user’s array aperture is negligible with respect to the propagation distance, we have
$\overline{\mathsf{v}}_{\varepsilon_{k}}\approx\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{s}}_{\varepsilon_{k}})\rvert^{2}A_{{\mathsf{u}},\varepsilon_{k}}^{2},{\mathsf{v}}_{\varepsilon_{k}}\approx\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{s}}_{\varepsilon_{k}})\rvert^{2}A_{{\mathsf{u}},\varepsilon_{k}},$ (45)
which simplifies $\gamma_{\mathsf{dl},\varepsilon_{k}}$ as follows:
$\gamma_{\mathsf{dl},\varepsilon_{k}}\approx\frac{A_{{\mathsf{u}},\varepsilon_{k}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{\varepsilon_{k}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{\varepsilon_{k}}^{2}+A_{{\mathsf{u}},\varepsilon_{k}}\sum_{k^{\prime}<k}{\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{\varepsilon_{k}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},\varepsilon_{k^{\prime}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}}.$ (46)
The sum-rate is given by $\sum_{k=1}^{K}\log_{2}(1+\gamma_{\mathsf{dl},\varepsilon_{k}})$.
Under the two-user case, without loss of generality, we consider the DPC order $2\rightarrow 1$. The achieved rates of user $1$ and user $2$ are given by
$\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}$ $=\log_{2}\left(1+\frac{A_{{\mathsf{u}},1}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{1}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{1}^{2}}\right),$ (47)
$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\frac{A_{{\mathsf{u}},2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{2}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{2}^{2}+A_{{\mathsf{u}},2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}_{2}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}\right),$ (48)
respectively. The problem of downlink sum-rate maximization is formulated as follows:
$\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}=\max\nolimits_{\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}}\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}+\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1},$ (49)
where $\mathsf{P}$ reflects the transmit power. Following the same derivation approach, the sum-rate achieved by the DPC order $1\rightarrow 2$ can also be derived. For Gaussian channels, the maximal DPC sum-rate is independent of the DPC decoding order, and thus $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}$ is essentially the downlink sum-rate capacity [21].
Problem (49) is non-convex, making the numerical identification of the maximum a nontrivial task. However, as demonstrated in [22], there exists a duality between the uplink and downlink scenarios. This duality establishes the equivalence of the DPC capacity region of a downlink channel with the capacity region of its dual uplink channel (as described in (24)), where users are subject to a sum power constraint. By defining $\overline{\gamma}_{\mathsf{du},k}\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}$ as the transmit SNRs of the dual uplink channel for $k=1,2$, the downlink sum-rate capacity is equal to the following dual uplink capacity [20, 21, 22]:
$\begin{split}{\mathsf{C}}_{\mathsf{du}}=\max_{{\mathsf{P}}_{1},{\mathsf{P}}_{2}\geq 0,{\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq{\mathsf{P}}}&\log_{2}(1+\overline{\gamma}_{\mathsf{du},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{2}\\ &+\overline{\gamma}_{\mathsf{du},1}\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}),\end{split}$ (50)
which is a convex problem and can be solved as follows.
Lemma 3.
The optimal power allocation for the dual uplink is given by
$({\mathsf{P}}_{1}^{\star},{\mathsf{P}}_{2}^{\star})=\begin{cases}({\mathsf{P}},0)&\xi\geq{\mathsf{P}}\\ (0,{\mathsf{P}})&\xi\leq-{\mathsf{P}}\\ (({{\mathsf{P}}-\xi})/{2},({{\mathsf{P}}+\xi})/{2})&\mathsf{else}\\ \end{cases},$ (51)
where $\xi=\frac{{A_{{\mathsf{u}},1}}\mathsf{g}_{1}/{\sigma_{1}^{2}}-{A_{{\mathsf{u}},2}}\mathsf{g}_{2}/{\sigma_{2}^{2}}}{{A_{{\mathsf{u}},1}A_{{\mathsf{u}},2}k_{0}^{2}\eta^{2}}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}/({4\pi\sigma_{1}^{2}\sigma_{2}^{2}})}$.
Proof:
The results can be readily obtained by utilizing the Karush–Kuhn–Tucker (KKT) conditions. ∎
Upon obtaining the optimal power allocation policy (51) for the dual uplink channel, we recover the corresponding downlink capacity-achieving source currents as follows.
Theorem 3.
Given the DPC order $2\rightarrow 1$ and the power constraint $\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}$, the downlink capacity-achieving source currents for user $1$ and user $2$ are given by
${\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})$ $=\sqrt{{\mathsf{P}}_{1}}\frac{\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})-\frac{{\mathsf{P}}_{2}\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})}{\sqrt{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}}},$ (52)
${\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})$ $=\sqrt{{\mathsf{P}}_{2}}\frac{\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})\sqrt{1+\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}},$ (53)
respectively, where $\hat{\mathsf{H}}_{k}(\mathbf{r})=\frac{\sqrt{A_{{\mathsf{u}},k}}}{\sigma_{k}}\mathsf{H}_{k}(\mathbf{r})$ for $k=1,2$. Additionally, the optimal source currents satisfy $\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}={\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$.
Proof:
Please refer to Appendix -D for more details. ∎
Remark 4.
Theorem 3 establishes the uplink-to-downlink transformation that takes as inputs an uplink transmit power allocation policy and an SIC decoding order, and outputs a set of downlink source currents with the same sum power as the uplink dual channel. As shown in Appendix -D and Section IV-B2, the output source currents achieve the same rates as the dual uplink channel using the input power allocation policy and SIC decoding with the specified decoding order.
Having obtained the capacity-achieving source currents under the DPC order $2\rightarrow 1$, we summarize the entire precoding procedure in Table II for ease of reference. The presented DPC procedure can be directly extended to the other order $1\rightarrow 2$ by exchanging the user indices.
1:Encode the data information into the symbols $\{s_{{\mathsf{dl}},k}\}_{k=1}^{2}$ using DPC with the order $2\rightarrow 1$; 2:Determine the dual power allocation policy $\{{\mathsf{P}}_{k}\}_{k=1}^{2}$ according to the specific target quality of service requirements; 3:Generate the source current ${\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})$: ${\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})=\sqrt{{\mathsf{P}}_{1}}\frac{\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})-\frac{{\mathsf{P}}_{2}\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})}{\sqrt{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}}};$ 4:Generate the source current ${\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})$: ${\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})=\sqrt{{\mathsf{P}}_{2}}\frac{\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})\sqrt{1+\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}};$ 5:Formulate the transmit signal ${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k}$.
TABLE II: DPC-based precoding for downlink CAPA communications.
IV-B2 Capacity Characterization
As proved in Appendix -D, given the power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, the achievable rates of user $1$ and user $2$ under the DPC order $2\rightarrow 1$ are given as follows:
$\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}\left(1-\frac{\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}|\rho|^{2}}{1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}}\right)\right),$ (54)
$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right).$ (55)
By comparing the above two equations with the results in Corollary 1, the following observations are made.
Remark 5.
The downlink user rates achieved by the DPC order $2\rightarrow 1$ equal those achieved in the dual uplink channel under the SIC decoding order $1\rightarrow 2$.
The achievable rates under the DPC order $1\rightarrow 2$ can be derived with similar steps, which are omitted for brevity. For a thorough study, we derive the downlink-to-uplink transformation as follows.
Theorem 4.
Given the source currents $\{{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\}_{k=1}^{2}$ such that $\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}$, the power allocation policy of the dual uplink channel, which achieves the same sum-rate as the downlink channel under the DPC order $2\rightarrow 1$, is given by
${\mathsf{P}}_{2}$ $=\frac{\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}(1+\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2})},$ (56)
${\mathsf{P}}_{1}$ $=\frac{\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}^{*}(\mathbf{r})\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}},$ (57)
respectively, where ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}$.
Proof:
Please refer to Appendix -E for more details. ∎
Remark 6.
Theorem 4 establishes the downlink-to-uplink transformation which, given a set of downlink source currents and a DPC encoding order, outputs the dual uplink power allocation policy with the same sum power as the downlink currents. This policy achieves uplink rates (using SIC decoding) equal to the downlink rates.
Remark 7.
The results in Remark 4 and Remark 6 suggest that under the same sum power constraint, the downlink DPC capacity region coincides with the capacity region of its dual uplink channel. This further verifies the existence of uplink-downlink duality in CAPA communications.
In light of the uplink-downlink duality, the sum-rate downlink capacity remains constant regardless of the DPC order (similar to its dual uplink channel), and is given as follows.
Theorem 5.
The sum-rate capacity of the downlink CAPA-based two-user channel is given by
$\mathsf{C}_{\mathsf{dl}}$ $=\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{dl},1}\overline{\gamma}_{\mathsf{dl},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}+\sum\nolimits_{k=1}^{2}\overline{\gamma}_{\mathsf{dl},k}{\mathsf{g}}_{k}\Big{)}$ (58a)
=$=\begin{cases}\log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},1}(\mathsf{P})\mathsf{g}_{1}\right)&\xi\geq{\mathsf{P}}\\ \log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},2}(\mathsf{P})\mathsf{g}_{2}\right)&\xi\leq-{\mathsf{P}}\\ \log_{2}\left(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{1}\varepsilon_{2}\overline{\rho}\right)&\mathrm{else}\\ \end{cases},$ (58b)
where $\widetilde{\gamma}_{\mathsf{dl},k}(x)\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}x$ for $k=1,2$, $\varepsilon_{1}=\widetilde{\gamma}_{\mathsf{dl},1}(\frac{{\mathsf{P}}-\xi}{2})\mathsf{g}_{1}$, and $\varepsilon_{2}=\widetilde{\gamma}_{\mathsf{dl},2}(\frac{{\mathsf{P}}+\xi}{2})\mathsf{g}_{2}$.
Proof:
Eq. (58) is obtained by inserting (51) into (50). ∎
Remark 8.
The downlink sum-rate capacity given in (58) is determined by the channel gain of each user and the channel correlation factor. This expression is applicable to an arbitrary aperture, regardless of its location, shape, and size.
Furthermore, according to the uplink-downlink duality, given a power allocation scheme $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ with ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, we can obtain the capacity region of a dual uplink channel. The downlink capacity region is the convex hull of the union of all these dual uplink capacity regions, which is illustrated in Section VI.
(a) Planar CAPA.
(b) Linear CAPA.
Figure 2: Geometries of CAPAs.
V Case Studies
As discussed in Remark 3 and Remark 8, the sum-rate capacity expressions derived for both uplink and downlink CAPA communications are applicable to apertures of arbitrary location, shape, and size, among other factors. In the following sections, we specialize ${\mathcal{A}}$ to several specific cases to unveil more system insights.
V-A Planar CAPAs
We first consider the case where the BS array is a planar CAPA placed on the $x$-$z$ plane and centered at the origin, as shown in Fig. 2(a). The edges of $\mathcal{A}$ are parallel to the axes, with physical dimensions $L_{x}$ and $L_{z}$ along the $x$- and $z$-axes, respectively. In this case, we have ${\mathcal{A}}=\{[x,0,z]^{\mathsf{T}}|x,z\in[-\frac{L_{x}}{2},\frac{L_{x}}{2}]\}$. For each user $k\in\{1,2\}$, let $r_{k}$ represent the distance from the center of $\mathcal{A}$ to the center of $\mathcal{A}_{\mathsf{k}}$, $\phi_{k}\in[0,\pi]$ and $\theta_{k}\in[0,\pi]$ denote the associated azimuth and elevation angles, respectively. Consequently, ${\mathbf{s}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Theta_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\cos{\phi_{k}}\sin{\theta_{k}}$, $\Psi_{k}\triangleq\sin{\phi_{k}}\sin{\theta_{k}}$, and $\Theta_{k}\triangleq\cos{\theta_{k}}$. Based on (8) and (15), for $\mathbf{r}=[x,0,z]^{\mathsf{T}}\in\mathcal{A}$, we have
$\begin{split}\mathsf{G}_{k}(\mathbf{r})&=\frac{\sqrt{r_{k}\Psi_{k}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{1}{2}}}}{\sqrt{4\pi}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{3}{4}}}\\ &\triangleq\mathsf{Q}_{k}(x,z).\end{split}$ (59)
To characterize the capacity, we derive closed-form expressions of the channel gain and correlation factor as follows.
Lemma 4.
With the planar CAPA, the channel gain for each user can be calculated as follows:
${\mathsf{g}}_{k}=\frac{1}{4\pi}\sum_{x\in{\mathcal{X}}_{k}}\sum_{z\in{\mathcal{Z}}_{k}}\arctan\bigg{(}\frac{xz/\Psi_{k}}{\sqrt{\Psi_{k}^{2}+x^{2}+z^{2}}}\bigg{)}\triangleq{\mathsf{g}}_{k}^{\mathsf{p}},$ (60)
where $k\in\{1,2\}$, ${\mathcal{X}}_{k}\triangleq\{\frac{L_{x}}{2r_{k}}\pm\Phi_{k}\}$, and ${\mathcal{Z}}_{k}\triangleq\{\frac{L_{z}}{2r_{k}}\pm\Theta_{k}\}$. The associated channel correlation factor is approximated by
$\begin{split}\rho=&\frac{\pi^{2}A}{4n^{2}\sqrt{\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}{\sqrt{\left(1-\psi_{j}^{2}\right)\left(1-\psi_{j^{\prime}}^{2}\right)}}\\ &\times\mathsf{Q}_{1}^{*}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\triangleq\rho_{\mathsf{p}},\end{split}$ (61)
where $n$ is a complexity-vs-accuracy tradeoff parameter, and $\psi_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$ for $j\in\{1,\ldots,n\}$.
Proof:
Please refer to Appendix -F for more details. ∎
Subsequently, the closed-form expressions of sum-rate capacity can be readily obtained.
Corollary 2.
The sum-rate capacity for the planar CAPA-based two-user channel can be written as follows:
$\mathsf{C}_{i}=\log_{2}\big{(}1+\overline{\gamma}_{i,1}\mathsf{g}_{1}^{\mathsf{p}}+\overline{\gamma}_{i,2}\mathsf{g}_{2}^{\mathsf{p}}+\overline{\gamma}_{i,1}\overline{\gamma}_{i,2}\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}\overline{\rho}_{\mathsf{p}}\big{)},$ (62)
where $i\in\{\mathsf{ul},\mathsf{dl}\}$, and $\overline{\rho}_{\mathsf{p}}=1-\lvert{\rho}_{\mathsf{p}}\rvert^{2}\in[0,1]$.
To shed more light on the characteristics of planar CAPAs, we investigate the asymptotic sum-rate capacity for a limiting case where $\mathcal{A}$ is assumed to be infinitely large, i.e., $L_{x},L_{z}\rightarrow\infty$. In this case, we have $\lim_{L_{x},L_{z}\rightarrow\infty}{\mathsf{g}}_{k}^{\mathsf{p}}=\frac{1}{4\pi}\frac{4\pi}{2}=\frac{1}{2}$. Regarding $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}$, while it is computationally intractable, based on the numerical results presented in [23, 24], we observe that $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}\ll 1$. Taken together, we can obtain the asymptotic sum-rate capacity as follows:
$\lim_{L_{x},L_{z}\rightarrow\infty}\mathsf{C}_{i}\approx\log_{2}(1+{\overline{\gamma}_{i,1}}/{2})+\log_{2}(1+{\overline{\gamma}_{i,2}}/{2}),$ (63)
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. The results in (63) suggest that when the BS is equipped with an infinitely large planar CAPA, half of the transmission power of each user can be captured by the receiver, and the inter-user interference vanishes.
Remark 9.
The above observation makes intuitive sense because an infinitely large planar CAPA can receive at most half of the power transmitted by an isotropic source, while the other half of the power will never reach the surface.
V-B Linear CAPAs
Next, we simplify the planar CAPA to a linear CAPA with $L_{x}\ll L_{z}$. In this case, the variations in currents and channel responses across the $x$-axis within $\mathcal{A}$ are negligible. The channel gain and correlation factor are calculated as follows.
Lemma 5.
The channel gain for the linear CAPA is given by
$\mathsf{g}_{k}=\frac{L_{x}\sin\phi_{k}\varrho_{k}}{4\pi r_{k}\sin\theta_{k}},$ (64)
where $\varrho_{k}=\frac{L_{z}-2r_{k}\Theta_{k}}{\left(L_{z}^{2}-4r_{k}\Theta_{k}L_{z}+4r_{k}^{2}\right)^{\frac{1}{2}}}+\frac{L_{z}+2r_{k}\Theta_{k}}{\left(L_{z}^{2}+4r_{k}\Theta_{k}L_{z}+4r_{k}^{2}\right)^{\frac{1}{2}}}$. The channel correlation factor satisfies
$\rho=\frac{\pi L_{z}}{2n}\sum_{j=1}^{n}{\sqrt{1-\psi_{j}^{2}}}\mathsf{Q}_{1}^{*}\Big{(}0,\frac{L_{z}\psi_{j}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}0,\frac{L_{z}\psi_{j}}{2}\Big{)}.$ (65)
Proof:
When $L_{x}\ll L_{z}$, we have $\mathsf{g}_{k}\approx L_{x}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\left|\mathsf{Q}_{k}(0,z)\right|^{2}\mathrm{d}z}$ and $\rho\approx L_{x}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\mathsf{Q}_{1}^{*}(0,z)\mathsf{Q}_{2}(0,z)\mathrm{d}z}$. The subsequent derivation steps are similar to those shown in Appendix -F, and the final results follow immediately. ∎
When the linear CAPA is infinitely long, i.e., $L_{z}\rightarrow\infty$, the asymptotic sum-rate capacity satisfies
$\lim_{L_{z}\rightarrow\infty}\mathsf{C}_{i}\approx\sum\nolimits_{k=1}^{2}\log_{2}\big{(}1+\frac{L_{x}\sin\phi_{1}{\overline{\gamma}_{i,k}}}{2\pi r_{k}\sin\theta_{k}}\big{)},$ (66)
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. By comparing (63) with (66), we observe that the asymptotic sum-rate capacity achieved by the linear CAPA is influenced not only by the transmit SNR but also by the locations of the users, which is in contrast to the planar CAPA.
Figure 3: Planar SPDA.
V-C Planar SPDAs
Finally, we consider a case where the above planar CAPA is partitioned into $M=M_{z}M_{x}$ spatially discrete elements, where $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$ denote the number of antenna elements along the $x$- and $z$-axes, as depicted in Fig. 3. The physical dimensions of each element are indicated by $\sqrt{{A}_{\mathsf{s}}}$, and the inter-element distance is denoted as $d$, where $\sqrt{{A}_{\mathsf{s}}}\leq d\ll r_{k}$. In this case, the central position of each element is given by $[m_{x}d,0,m_{z}d]^{\mathsf{T}}$ for $m_{x}\in\mathcal{M}_{x}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{x}\}$ and $m_{z}\in\mathcal{M}_{z}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{z}\}$. Further, we have $L_{x}\approx M_{x}d$, $L_{z}\approx M_{z}d$, and $\mathcal{A}=\{[m_{x}d+\ell_{x},0,m_{z}d+\ell_{z}]^{\mathsf{T}}|\ell_{x},\ell_{z}\in[-\frac{\sqrt{{A}_{\mathsf{s}}}}{2},\frac{\sqrt{{A}_{\mathsf{s}}}}{2}],m_{x}\in\mathcal{M}_{x},m_{z}\in\mathcal{M}_{z}\}$.
Lemma 6.
The channel gain and channel correlation factor for the planar SPDA satisfy ${\mathsf{g}}_{k}^{\mathsf{s}}=\zeta_{\mathsf{oc}}{\mathsf{g}}_{k}^{\mathsf{p}}$ and $\rho_{\mathsf{s}}={A}_{\mathsf{s}}\sum_{m_{x}\in\mathcal{M}_{x}}{\sum_{m_{z}\in\mathcal{M}_{z}}{\mathsf{Q}_{1}^{*}(m_{x}d,m_{z}d)\mathsf{Q}_{2}(m_{x}d,m_{z}d)}}$, respectively, where $\zeta_{\mathsf{oc}}\triangleq\frac{{A}_{\mathsf{s}}}{d^{2}}\in(0,1]$ is defined as the array occupation ratio.
Proof:
Please refer to Appendix -G for more details. ∎
Therefore, the asymptotic sum-rate capacity achieved by the planar SPDA is given by
$\lim_{M_{x},M_{z}\rightarrow\infty}\mathsf{C}_{i}\approx\sum\nolimits_{k=1}^{2}\log_{2}(1+\zeta_{\mathsf{oc}}{\overline{\gamma}_{i,k}}/{2}),$ (67)
where $i\in\{{\mathsf{ul}},{\mathsf{dl}}\}$. Comparing (63) with (67) yields the following observations.
Remark 10.
The capacity achieved by an SPDA increases to that achieved by a CAPA when $\zeta_{\mathsf{oc}}=1$. This makes intuitive sense as an SPDA turns into a CAPA when the array occupation ratio equals one.
VI Numerical Results
In this section, computer simulations are performed to demonstrate the performance of CAPA communications and verify the accuracy of the developed analytical results. For clarity, the simulations employ the planar arrays specified in Section V. Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.0107$ m, $d=\frac{\lambda}{2}$, $A_{\mathsf{u},1}=A_{\mathsf{u},2}=\frac{\lambda^{2}}{4\pi}$, $L_{x}=L_{z}$, $M_{x}=M_{z}$, $\overline{\gamma}_{\mathsf{ul},1}=30$ dB, $\overline{\gamma}_{\mathsf{ul},2}=40$ dB, $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}=50$ dB with $\sigma_{1}^{2}=\sigma_{2}^{2}$, $\theta_{1}=\theta_{2}=\frac{\pi}{6}$, $\phi_{1}=\phi_{2}=\frac{\pi}{3}$, $r_{1}=10$ m, and $r_{2}=20$ m. All the results are calculated by setting the complexity-vs-accuracy tradeoff parameter $n$ as $n=200$.
Figure 4: Uplink transmission rates achieved by CAPA.
(a) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
(b) Sum-rate capacity vs. $A$.
Figure 5: Uplink sum-rate capacity.
VI-A Uplink CAPA Communications
Figure 4 illustrates the uplink per-user rate and sum-rate capacity achieved by the CAPA in terms of the BS aperture size under the SIC decoding orders $1\rightarrow 2$ and $2\rightarrow 1$. It can be observed from this graph that as the aperture size $A$ grows, all the presented rates increase. This increase is attributed to the larger aperture size enhancing channel gain and reducing channel correlation, which in turn enhances the dedicated signal and diminishes the IUI, as discussed in (63).
To compare the performance gap between CAPA and the conventional SPDA in terms of sum-rate capacity, we plot the sum-rate capacity for SPDAs as a function of the array occupation ratio $\zeta_{\mathsf{oc}}$ in Figure 5(a). As shown in the graph, the sum-rate capacity for an SPDA gradually converges with that of a CAPA as the array occupation ratio increases towards $1$. This observation validates the statements in Remark 10. Additionally, the results in Figure 5(a) indicate that increasing the aperture size can boost the channel capacity. To gain further insights into the capacity limit as the BS aperture size $A$ approaches infinity, we plot the sum-rate capacity against the aperture size in Figure 5(b). As anticipated, increasing the aperture size causes the sum-rate capacity to approach its upper bound, as derived in (63) and (67). It is worth mentioning that the asymptotic sum-rate capacity for an infinitely large aperture is a finite value, subject to the principle of energy conservation [23]. This verifies our discussions in Remark 9.
(a) Capacity region vs. $A$
(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$.
Figure 6: Uplink capacity regions.
Figure 6 shows the uplink capacity region achieved by the CAPA, where the green and blue points represent the achieved rates by SIC decoding order $1\rightarrow 2$ and $2\rightarrow 1$, respectively. The rate tuple on the line segment connecting these two points is achieved by the time-sharing strategy. In Figure 6(a), the capacity regions achieved by the CAPA for various aperture sizes are illustrated. It can be observed that the capacity region expands with increasing $A$, shifting from a pentagonal to a rectangular shape. This phenomenon occurs because the IUI decreases as the aperture size increases, which is consistent with the results shown in Figure 4. Furthermore, as depicted in Figure 6(b), the capacity region of the CAPA surpasses that of the SPDA, and their disparity narrows as the array occupation rate $\zeta_{\mathsf{oc}}$ increases, reinforcing findings from Figure 5(a). These findings underscore the superiority of CAPAs over conventional SPDAs in terms of channel capacity.
Figure 7: Downlink transmission rates achieved by CAPA.
(a) Sum-rate capacity vs. $A$.
(b) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
Figure 8: Downlink sum-rate capacity.
VI-B Downlink CAPA Communications
Figure 7 plots the per-user rate and the sum-rate capacity in terms of the total transmit SNR $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}$ in downlink CAPA communications. It is observed that the transmission rates of both users exhibit identical high-SNR slopes, regardless of the DPC encoding orders. Figure 8(a) and Figure 8(b) illustrate the downlink sum-rate capacity with respect to the aperture size and array occupation ratio, respectively. Similar to the uplink case, the downlink sum-rate capacity converges to its upper bound as the aperture size increases. Additionally, for the SPDA, the achieved sum-rate capacity is positively correlated with the array occupation ratio, further suggesting that the CAPA outperforms the SPDA in terms of channel capacity.
(a) CAPA capacity region vs. $A$.
(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$
Figure 9: Downlink capacity regions.
Figure 9 illustrates the downlink capacity region achieved by the CAPA. The downlink capacity region is depicted by the convex hull of all the corresponding dual uplink capacity regions with the same sum power budget [20]. This means that the boundary of the downlink capacity region is formed by the corner points of the dual uplink capacity regions, as illustrated in Figure 9(a). It is observed that the downlink capacity region of the CAPA expands with the aperture size $A$, whereas the capacity region of the SPDA expands with the array occupation ratio $\zeta_{\mathsf{oc}}$. Taken together, these observations demonstrate that the CAPA achieves a more extensive downlink capacity region than the conventional SPDA, highlighting its superiority.
VII Conclusion
This article has proposed an analytically tractable framework for CAPA-based wireless communications by utilizing continuous operators to address the continuous nature of the EM field. This framework significantly differs from the conventional matrix-based transmission framework for SPDA-based communications. Based on this novel framework, we have characterized the capacity limits of uplink and downlink channels for CAPAs in both single-user and multiuser cases. For each investigated scenario, we derived closed-form expressions for the capacity-achieving detectors or source currents along with the corresponding capacity. Additionally, we have analyzed the channel capacity by specializing the derived results to several typical array structures. Through both theoretical analyses and numerical simulations, we have demonstrated that CAPA exhibits a higher sum-rate capacity and a more extensive capacity region compared to the conventional SPDA. This underscores CAPA’s potential as a promising transmission paradigm for future wireless networks.
-A Proof of Lemma 1
The mean of $\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}$ is calculated as follows:
${\mathbbmss{E}}\left\{\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\right\}=\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathbbmss{E}}\{{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})\}{\rm{d}}{\mathbf{r}},$ (A1)
which, together with the fact that ${\mathbbmss{E}}\{{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r})\}=0$, yields ${\mathbbmss{E}}\left\{\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\right\}=0$. The variance is given by
${\mathbbmss{E}}\left\{\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r}^{\prime}){\mathsf{N}}_{\mathsf{ul}}^{*}(\mathbf{r}^{\prime}){\rm{d}}{\mathbf{r}}^{\prime}\right\}$ (A2)
$=\int_{{\mathcal{A}}}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{V}}_{\mathsf{ul}}(\mathbf{r}^{\prime}){\mathbbmss{E}}\{{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}^{*}(\mathbf{r}^{\prime})\}{\rm{d}}{\mathbf{r}}{\rm{d}}{\mathbf{r}}^{\prime}$ (A3)
$=\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r}^{\prime})\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r})\sigma^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime}){\rm{d}}{\mathbf{r}}{\rm{d}}{\mathbf{r}}^{\prime}$ (A4)
$\overset{\clubsuit}{=}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r}^{\prime})(\sigma^{2}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}^{\prime})){\rm{d}}{\mathbf{r}}^{\prime}=\sigma^{2}\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}},$ (A5)
where the step $\clubsuit$ is based on the fact that $\int_{{\mathcal{A}}}\delta({\mathbf{x}}-{\mathbf{x}}_{0})f(\mathbf{x}){\rm{d}}{\mathbf{x}}=f({\mathbf{x}}_{0})$ with $f(\cdot)$ being an arbitrary function defined on ${\mathcal{A}}$.
-B Proof of Lemma 2
Firstly, we prove that the autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r})$ satisfies ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$. Specifically, ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}$ can be calculated as follows:
${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}$
$=\int_{{\mathcal{A}}}\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{x}){\mathsf{W}}_{{\mathsf{Z}}}^{*}(\mathbf{r}^{\prime},\mathbf{x}^{\prime}){\mathbbmss{E}}\{{\mathsf{Z}}(\mathbf{x}){\mathsf{Z}}^{*}(\mathbf{x}^{\prime})\}{\rm{d}}\mathbf{x}{\rm{d}}\mathbf{x}^{\prime}$ (A6)
$=\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}^{*}(\mathbf{r}^{\prime},\mathbf{x}^{\prime})\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{x}){\mathsf{R}}_{{\mathsf{Z}}}(\mathbf{x},{\mathbf{x}}^{\prime}){\rm{d}}\mathbf{x}{\rm{d}}\mathbf{x}^{\prime}.$ (A7)
By substituting (27) and (30) into (A7) and then calculating the inner integral with respect to $\mathbf{x}$, we obtain
$\begin{split}&\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{x}){\mathsf{R}}_{{\mathsf{Z}}}(\mathbf{x},{\mathbf{x}}^{\prime}){\rm{d}}\mathbf{x}=\sigma^{2}(\delta({\mathbf{r}}-{\mathbf{x}}^{\prime})\\ &+(\overline{\gamma}_{\mathsf{ul},1}+\mu_{1}+\mu_{1}\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}){\mathsf{G}}_{1}(\mathbf{r}){\mathsf{G}}_{1}^{*}(\mathbf{x}^{\prime})).\end{split}$ (A8)
By continuously substituting (A8) and (30) into (A7), we next calculate the outer integral in terms of ${{\mathbf{x}}}^{\prime}$, which yields
$\begin{split}&{\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}=\sigma^{2}(\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})+{\mathsf{G}}_{1}(\mathbf{r}){\mathsf{G}}_{1}^{*}(\mathbf{r}^{\prime})\\ &\times(2\mu_{1}+\overline{\gamma}_{\mathsf{ul},1}+2\mu_{1}\overline{\gamma}_{\mathsf{ul},1}{{\mathsf{g}}_{1}}+\mu_{1}^{2}{{\mathsf{g}}_{1}}+\overline{\gamma}_{\mathsf{ul},1}\mu_{1}^{2}{{\mathsf{g}}_{1}^{2}})).\end{split}$ (A9)
With $\mu_{1}=-\frac{1}{{\mathsf{g}}_{1}}\pm\frac{1}{{\mathsf{g}}_{1}\sqrt{1+\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}}}$, it is readily shown that
$0\displaystyle 2\mu_{1}+\overline{\gamma}_{\mathsf{ul},1}+2\mu_{1}\overline{\gamma}_{\mathsf{ul},1}{{\mathsf{g}}_{1}}+\mu_{1}^{2}{{\mathsf{g}}_{1}}+\overline{\gamma}_{\mathsf{ul},1}\mu_{1}^{2}{{\mathsf{g}}_{1}^{2}}=$ 0 . (A10)
Therefore, we have ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$.
Next, we prove that ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ is an invertible linear transformation. We define
$\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{r}^{\prime})\triangleq\delta(\mathbf{r}-\mathbf{r}^{\prime})-\frac{\mu_{1}}{1+\mu_{1}{\mathsf{g}}_{1}}{\mathsf{G}}_{1}(\mathbf{r}){\mathsf{G}}_{1}^{*}(\mathbf{r}^{\prime}).$ (A11)
For an arbitrary function ${f}(\mathbf{r})$ defined in $\mathbf{r}\in{\mathcal{A}}$, we have
$\begin{split}&\int_{{\mathcal{A}}}\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime})\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})f(\mathbf{r}){\rm{d}}\mathbf{r}{\rm{d}}\mathbf{r}^{\prime}\\ &=\int_{{\mathcal{A}}}{f}(\mathbf{r})\int_{{\mathcal{A}}}\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime}){\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\rm{d}}\mathbf{r}^{\prime}{\rm{d}}\mathbf{r}.\end{split}$ (A12)
By inserting (30) and (A11) into (A12) and utilizing the fact that $\int_{{\mathcal{A}}}\delta({\mathbf{x}}-{\mathbf{x}}_{0})f(\mathbf{x}){\rm{d}}{\mathbf{x}}=f({\mathbf{x}}_{0})$, we calculate the inner integral in terms of $\mathbf{r}^{\prime}$ as follows:
$\begin{split}&\int_{{\mathcal{A}}}\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime}){\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\rm{d}}\mathbf{r}^{\prime}=\delta({\mathbf{x}}-{\mathbf{r}})+{\mathsf{G}}_{1}(\mathbf{x})\\ &\times{\mathsf{G}}_{1}^{*}(\mathbf{r})\bigg{(}\mu_{1}-\frac{\mu_{1}+\mu_{1}^{2}{\mathsf{g}}_{1}}{1+\mu_{1}{\mathsf{g}}_{1}}\bigg{)}=\delta({\mathbf{x}}-{\mathbf{r}}).\end{split}$ (A13)
Plugging (A13) into (A12) gives
$\begin{split}&\int_{{\mathcal{A}}}\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{x},\mathbf{r}^{\prime})\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})f(\mathbf{r}){\rm{d}}\mathbf{r}{\rm{d}}\mathbf{r}^{\prime}\\ &=\int_{{\mathcal{A}}}{f}(\mathbf{r})\delta({\mathbf{x}}-{\mathbf{r}}){\rm{d}}\mathbf{r}={f}(\mathbf{x}),\end{split}$ (A14)
which suggests that the inversion of ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ is $\overline{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r},\mathbf{r}^{\prime})$.
-C Proof of Theorem 1
By definition, $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ can be calculated as follows:
$\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ $=\int_{{\mathcal{A}}}(\delta(\mathbf{r}^{\prime}-\mathbf{r})+\mu_{1}{\mathsf{G}}_{1}(\mathbf{r}^{\prime}){\mathsf{G}}_{1}^{*}(\mathbf{r})){\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}\mathbf{r}$ (A15)
$={{\rm{j}}k_{0}\eta}/{\sqrt{4\pi}}({\mathsf{G}}_{2}(\mathbf{r}^{\prime})+\mu_{1}\sqrt{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}\rho{\mathsf{G}}_{1}(\mathbf{r}^{\prime})).$ (A16)
Inserting (A16) into (33) yields
$\gamma_{\mathsf{ul},2}$ $=\overline{\gamma}_{\mathsf{ul},2}\int_{{\mathcal{A}}}\lvert{\mathsf{G}}_{2}(\mathbf{r}^{\prime})+\mu_{1}\sqrt{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}\rho{\mathsf{G}}_{1}(\mathbf{r}^{\prime})\rvert^{2}{\rm{d}}{\mathbf{r}}^{\prime}$ (A17)
$=\overline{\gamma}_{\mathsf{ul},2}\int_{{\mathcal{A}}}\lvert{\mathsf{G}}_{2}(\mathbf{r}^{\prime})\rvert^{2}{\rm{d}}{\mathbf{r}}^{\prime}+\overline{\gamma}_{\mathsf{ul},2}{\mu_{1}^{2}}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\lvert\rho\rvert^{2}\int_{{\mathcal{A}}}\lvert{\mathsf{G}}_{1}(\mathbf{r}^{\prime})\rvert^{2}{\rm{d}}{\mathbf{r}}^{\prime}$
$+2\overline{\gamma}_{\mathsf{ul},2}\mu_{1}\sqrt{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}\Re\left\{\rho\int_{{\mathcal{A}}}{\mathsf{G}}_{1}({\mathbf{r}}){\mathsf{G}}_{2}^{*}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right\}$ (A18)
$=\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{2}(1+{\mathsf{g}}_{1}\lvert\rho\rvert^{2}\mu_{1}({\mu_{1}}{\mathsf{g}}_{1}+2)).$ (A19)
Furthermore, based on (A10), we have
$\begin{split}\mu_{1}({\mu_{1}}{{\mathsf{g}}_{1}}+2)=-\overline{\gamma}_{\mathsf{ul},1}(1+\mu_{1}{\mathsf{g}}_{1})^{2},\end{split}$ (A20)
which, together with (A19), yields
$\gamma_{\mathsf{ul},2}=\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{2}(1-\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}\lvert\rho\rvert^{2}(1+\mu_{1}{\mathsf{g}}_{1})^{2}).$ (A21)
Plugging $\mu_{1}=-\frac{1}{{\mathsf{g}}_{1}}\pm\frac{1}{{\mathsf{g}}_{1}\sqrt{1+\overline{\gamma}_{\mathsf{ul},1}{\mathsf{g}}_{1}}}$ into (A21) gives
$\gamma_{\mathsf{ul},2}=\overline{\gamma}_{\mathsf{ul},2}\mathsf{g}_{2}\left(1-\frac{\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}\left|\rho\right|^{2}}{1+\overline{\gamma}_{\mathsf{ul},1}\mathsf{g}_{1}}\right).$ (A22)
The final results follow immediately.
-D Proof of Theorem 3
Our objective is to verify that the presented source current distributions $\{{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\}_{k=1}^{2}$ are able to: i) achieve the downlink sum-rate capacity $\log_{2}(1+\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}+\overline{\gamma}_{\mathsf{dl},1}\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho})$, and ii) satisfy the sum power constraint $\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}={\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$.
Substituting (52) into $\lvert\int_{{\mathcal{A}}}\hat{\mathsf{H}}_{1}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}$ gives
$\begin{split}&\left\lvert\int_{{\mathcal{A}}}\hat{\mathsf{H}}_{1}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right\rvert^{2}=\frac{{{\mathsf{P}}_{1}}{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{-1-{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\\ &+{{\mathsf{P}}_{1}}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}=\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}\left(1-\frac{\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}|\rho|^{2}}{1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}}\right),\end{split}$ (A23)
which yields
$\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}$ $=\log_{2}\left(1+\left\lvert\int_{{\mathcal{A}}}\hat{\mathsf{H}}_{1}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right\rvert^{2}\right)$ (A24)
$=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}\left(1-\frac{\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}|\rho|^{2}}{1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}}\right)\right).$ (A25)
Moreover, inserting (53) into (48) gives
$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}=\log_{2}\left(1+\frac{\lvert\int_{{\mathcal{A}}}\hat{\mathsf{H}}_{2}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+\lvert\int_{{\mathcal{A}}}\hat{\mathsf{H}}_{2}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}\right)$ (A26)
$=\log_{2}\left(1+{\mathsf{P}}_{2}\frac{\lvert\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}\rvert^{2}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\right)$ (A27)
$=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right).$ (A28)
Inserting (A25) and (A28) into $\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}+\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ gives the sum-rate capacity for a given $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, i.e., $\log_{2}(1+\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}+\overline{\gamma}_{\mathsf{dl},1}\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho})$, which concludes the proof of the first part.
We next turn to the sum power constraint. Note that
$\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},1}(\mathbf{r})\rvert^{2}}\mathrm{d}\mathbf{r}=\int_{\mathcal{A}}{\int_{\mathcal{A}}{\mathsf{J}}_{{\mathsf{dl}},1}(\mathbf{r})\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right){\mathsf{J}}_{{\mathsf{dl}},1}^{*}(\mathbf{r}^{\prime})}\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{r}^{\prime}.$ (A29)
On the other hand, based on (53), $\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},2}(\mathbf{r})\rvert^{2}}\mathrm{d}\mathbf{r}$ can be calculated as follows:
$\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},2}(\mathbf{r})\rvert^{2}}\mathrm{d}\mathbf{r}={{\mathsf{P}}_{2}}+{{\mathsf{P}}_{2}}\left\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right\rvert^{2}$ (A30)
$={{\mathsf{P}}_{2}}+{{\mathsf{P}}_{2}}\int_{\mathcal{A}}\int_{\mathcal{A}}{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r})\hat{\mathsf{\mathsf{H}}}_{2}^{*}(\mathbf{r}^{\prime}){\mathsf{J}}_{{\mathsf{dl}},1}^{*}({\mathbf{r}}){\rm{d}}{\mathbf{r}}{\rm{d}}{\mathbf{r}}.$ (A31)
The sum of (A29) and (A31) satisfies
$\begin{split}\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}&={{\mathsf{P}}_{2}}+\int_{\mathcal{A}}\int_{\mathcal{A}}{\mathsf{J}_{{\mathsf{dl}},1}(\mathbf{r})}(\delta(\mathbf{r}-\mathbf{r}^{\prime})\\ &+{{\mathsf{P}}_{2}}{\hat{\mathsf{H}}}_{2}(\mathbf{r})\hat{\mathsf{H}}_{2}^{*}(\mathbf{r}^{\prime}))\mathrm{d}\mathbf{r}{\mathsf{J}}_{{\mathsf{dl}},1}^{*}(\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}.\end{split}$ (A32)
Substituting (52) into (A32) yields
$\int_{\mathcal{A}}\int_{\mathcal{A}}{\mathsf{J}_{{\mathsf{dl}},1}(\mathbf{r})}(\delta(\mathbf{r}-\mathbf{r}^{\prime})+{{\mathsf{P}}_{2}}{\hat{\mathsf{H}}}_{2}(\mathbf{r})\hat{\mathsf{H}}_{2}^{*}(\mathbf{r}^{\prime}))\mathrm{d}\mathbf{r}{\mathsf{J}}_{{\mathsf{dl}},1}^{*}(\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}$
$={\mathsf{P}}_{1}\frac{\int_{\mathcal{A}}\int_{\mathcal{A}}f_{\mathsf{n}}(\mathbf{r})(\delta(\mathbf{r}-\mathbf{r}^{\prime})+{{\mathsf{P}}_{2}}{\hat{\mathsf{H}}}_{2}(\mathbf{r})\hat{\mathsf{H}}_{2}^{*}(\mathbf{r}^{\prime}))\mathrm{d}\mathbf{r}f_{\mathsf{n}}^{*}(\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}},$ (A33)
where $f_{\mathsf{n}}(\mathbf{r})\triangleq\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})-\frac{{\mathsf{P}}_{2}\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})$. The inner integral in the numerator of (A33) can be calculated as follows:
$\int_{\mathcal{A}}f_{\mathsf{n}}(\mathbf{r})(\delta(\mathbf{r}-\mathbf{r}^{\prime})+{{\mathsf{P}}_{2}}{\hat{\mathsf{H}}}_{2}(\mathbf{r})\hat{\mathsf{H}}_{2}^{*}(\mathbf{r}^{\prime}))\mathrm{d}\mathbf{r}$ (A34)
$=\hat{\mathsf{H}}_{1}^{*}(\mathbf{r}^{\prime})+\bigg{(}1-\frac{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\bigg{)}$
×$\times{\mathsf{P}}_{2}\bigg{(}\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\bigg{)}\hat{\mathsf{H}}_{2}^{*}(\mathbf{r}^{\prime})=\hat{\mathsf{H}}_{1}^{*}(\mathbf{r}^{\prime}),$ (A35)
based on which the outer integral in the numerator of (A33) is calculated as follows:
$\begin{split}\int_{\mathcal{A}}\!\hat{\mathsf{H}}_{1}^{*}(\mathbf{r}^{\prime})f_{\mathsf{n}}^{*}(\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}\!\!=\!-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\!\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\!+\!\!\int_{\mathcal{A}}\!\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}.\end{split}$ (A36)
As a result, we obtain
(A33)$={\mathsf{P}}_{1}\frac{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}}={\mathsf{P}}_{1},$ (A37)
which, together with (A32), suggests that $\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}={\mathsf{P}}_{1}+{\mathsf{P}}_{2}$. As the downlink sum-rate equals that of its dual uplink channel under the same sum power constraint, the downlink capacity is achieved when the CAPA transmits at maximum power, i.e., ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, which concludes the proof of the second part.
-E Proof of Theorem 4
Our objective is to verify that the presented power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ is able to: i) ensure the uplink rates equal the downlink rates shown in (47) and (48), and ii) satisfy the sum power constraint ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}$. The first part can be directly proved by substituting (56) and (57) into (54) and (55). We next turn to the sum power constraint.
Following the derivation steps in Appendix -B, it is proved that the inversion of ${\mathsf{U}}(\mathbf{r}^{\prime},\mathbf{r})$ is $\overline{\mathsf{U}}(\mathbf{r},\mathbf{r}^{\prime})$, where ${\mathsf{U}}(\mathbf{r}^{\prime},\mathbf{r})=\delta(\mathbf{r}^{\prime}-\mathbf{r})+\mu_{2}{\mathsf{G}}_{2}(\mathbf{r}^{\prime}){\mathsf{G}}_{2}^{*}(\mathbf{r})$ with $\mu_{2}=-\frac{1}{{\mathsf{g}}_{2}}\pm\frac{1}{{\mathsf{g}}_{2}\sqrt{1+\overline{\gamma}_{\mathsf{ul},2}{\mathsf{g}}_{2}}}$ and $\overline{\mathsf{U}}(\mathbf{r},\mathbf{r}^{\prime})=\delta(\mathbf{r}-\mathbf{r}^{\prime})-\frac{\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}}{\mathsf{G}}_{2}(\mathbf{r}){\mathsf{G}}_{2}^{*}(\mathbf{r}^{\prime})$, i.e., $\int_{\mathcal{A}}\overline{\mathsf{U}}(\mathbf{r},{\mathbf{x}}){\mathsf{U}}(\mathbf{x},\mathbf{r}^{\prime}){\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}{\mathsf{U}}(\mathbf{r},{\mathbf{x}})\overline{\mathsf{U}}(\mathbf{x},\mathbf{r}^{\prime}){\rm{d}}{\mathbf{x}}=\delta({\mathbf{r}}-{\mathbf{r}}^{\prime})$. Furthermore, it is worth noting that $\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}=\int_{\mathcal{A}}\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}^{\prime})\delta(\mathbf{r}^{\prime}-\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}^{\prime}{\rm{d}}{\mathbf{r}}\triangleq{\diamond}$. It follows that
[…]
$=\int_{\mathcal{A}}\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}^{\prime}){\mathsf{U}}(\mathbf{r}^{\prime},{\mathbf{x}}){\rm{d}}{\mathbf{r}}^{\prime}\int_{\mathcal{A}}\overline{\mathsf{U}}(\mathbf{x},\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}{\rm{d}}{\mathbf{x}},$ (A38)
which, together with the Cauchy’s inequality, yields
$\lvert{\diamond}\rvert^{2}\leq\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}\int_{\mathcal{A}}\lvert{\diamond}_{2}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}},$ (A39)
where ${\diamond}_{1}(\mathbf{x})\triangleq\int_{\mathcal{A}}\overline{\mathsf{U}}(\mathbf{x},\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}={\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{x}})-\frac{\mu_{2}{\mathsf{G}}_{2}^{*}(\mathbf{x})}{1+\mu_{2}{\mathsf{g}}_{2}}\int_{\mathcal{A}}{\mathsf{G}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}$ and ${\diamond}_{2}(\mathbf{x})\triangleq\int_{\mathcal{A}}\hat{{\mathsf{H}}}_{1}(\mathbf{r}^{\prime}){\mathsf{U}}(\mathbf{r}^{\prime},{\mathbf{x}}){\rm{d}}{\mathbf{r}}^{\prime}=\mu_{2}{\mathsf{G}}_{2}(\mathbf{x})\int_{\mathcal{A}}{\mathsf{G}}_{2}^{*}(\mathbf{r}^{\prime})\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}^{\prime}){\rm{d}}{\mathbf{r}}^{\prime}+\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{x})$. Consequently, we have
$\int_{\mathcal{A}}\lvert{\diamond}_{2}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert\hat{{\mathsf{H}}}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}+\Big{\lvert}\int_{\mathcal{A}}{\mathsf{G}}_{2}^{*}(\mathbf{r}^{\prime})\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r}^{\prime}){\rm{d}}{\mathbf{r}}^{\prime}\Big{\rvert}^{2}$
×$\times(2\mu_{2}+\mu_{2}^{2}{\mathsf{g}}_{2})=\int_{\mathcal{A}}\lvert\hat{{\mathsf{H}}}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}-\frac{\lvert\int_{\mathcal{A}}{\mathsf{G}}_{2}(\mathbf{r}^{\prime})\hat{\mathsf{\mathsf{H}}}_{1}^{*}(\mathbf{r}^{\prime}){\rm{d}}{\mathbf{r}}^{\prime}\rvert^{2}}{(1+\overline{\gamma}_{\mathsf{dl},2}{\mathsf{g}}_{2})/\overline{\gamma}_{\mathsf{dl},2}}$
$=\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{1}^{*}(\mathbf{r})\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}},$ (A40)
which, together with (57) and (A39), yields ${\mathsf{P}}_{1}\leq\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}$. Furthermore, we obtain
$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}-\Big{\lvert}\int_{\mathcal{A}}{\mathsf{G}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}$
×$\times\Big{(}\frac{2\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}}-\Big{(}\frac{\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}}\Big{)}^{2}{\mathsf{g}}_{2}\Big{)}.$ (A41)
Recalling that $\mu_{2}=-\frac{1}{{\mathsf{g}}_{2}}\pm\frac{1}{{\mathsf{g}}_{2}\sqrt{1+\overline{\gamma}_{\mathsf{dl},2}{\mathsf{g}}_{2}}}$ yields $\frac{2\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}}-(\frac{\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}})^{2}{\mathsf{g}}_{2}=-\overline{\gamma}_{\mathsf{dl},2}$, which yields
$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}.$
As proved above, we have ${\mathsf{P}}_{1}\leq\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}$, and thus the powers satisfy
${\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{(}1+\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}\Big{)},$ (A42)
which, together with (56), yields
${\mathsf{P}}_{1}+{\mathsf{P}}_{2}$ $\leq\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+\frac{\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}},$
$\overset{\heartsuit}{\leq}\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}\leq{\mathsf{P}},$ (A43)
where the step $\heartsuit$ in (A43) is due to the Cauchy’s inequality. This completes the proof of Theorem 4.
-F Proof of Lemma 4
Based on (59), the channel gain for the planar CAPA can be calculated as follows:
$\mathsf{g}_{k}=\int_{\frac{L_{z}}{-2}}^{\frac{L_{z}}{2}}{\int_{\frac{L_{x}}{-2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{k}^{*}(x,z)\mathsf{Q}_{k}(x,z)\mathrm{d}x}\mathrm{d}z}=\frac{r_{k}\Psi_{k}}{4\pi}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}$
×$\times{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}}{{(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{-\frac{3}{2}}\mathrm{d}x}\mathrm{d}z}.$
Subsequently, we can calculate the inner integral with the aid of [25, Eq. (2.264.5)] and the outer integral with the aid of [25, Eq. (2.284)], which yields the results in (60). The channel correlation factor can be written as follows:
$\rho=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{1}^{*}(x,z)\mathsf{Q}_{2}(x,z)\mathrm{d}x}\mathrm{d}z}.$ (A44)
The above integrals can be calculated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\sum_{j=1}^{n}{f\left(x_{j}\right)}$ with $x_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$, lead to the results in (61). This completes the proof of Lemma 4.
-G Proof of Lemma 6
In the case of an planar SPDA, due to the small element size compared to the distance between the BS and the user, i.e., $\sqrt{{A}_{\mathsf{s}}}\ll r_{k}$, the variation of the channel across an element is negligible. Thus, the channel gain can be written as follows:
$\mathsf{g}_{k}=\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}{\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}{{{A}_{\mathsf{s}}}\left|\mathsf{Q}_{k}(m_{x}d,m_{z}d)\right|}}^{2}.$ (A45)
Defining $\epsilon_{k}=\frac{d}{r_{k}}\ll 1$, we can rewrite (A45) as follows:
$\begin{split}\mathsf{g}_{k}=\frac{A\Psi_{k}}{4\pi r_{k}^{2}}\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right),\end{split}$ (A46)
where $f_{\mathsf{sp}}^{(k)}\left(x,z\right)\triangleq(x^{2}+z^{2}-2\Phi_{k}x-2\Omega_{k}z+1)^{-\frac{3}{2}}$ is defined within the square area ${\mathcal{C}}_{k}\triangleq\{(x,z)\mid-\frac{M_{x}\epsilon_{k}}{2}\leq x\leq\frac{M_{x}\epsilon_{k}}{2},-\frac{M_{z}\epsilon_{k}}{2}\leq z\leq\frac{M_{z}\epsilon_{k}}{2}\}$, which is then divided into $M_{x}M_{z}$ sub-squares, each with an area $\epsilon_{k}^{2}$. As $\epsilon_{k}\ll 1$, we have $f_{\mathsf{sp}}^{(k)}\left(x,z\right)\approx f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)$ for $\forall\left(x,z\right)\in\left\{\left(x,z\right)\mid\left(m_{x}-\frac{1}{2}\right)\epsilon_{k}\leq x\leq\left(m_{x}+\frac{1}{2}\right)\epsilon_{k},\left(m_{z}-\frac{1}{2}\right)\epsilon_{k}\leq z\right.$ $\left.\leq\left(m_{z}+\frac{1}{2}\right)\epsilon_{k}\right\}$. It follows from the concept of double integral that
$\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}\!\!\!{f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)\epsilon_{k}^{2}}\approx\!\!\iint_{{\mathcal{C}}_{k}}{f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$
Consequently, (A46) can be rewritten as follows:
$\begin{split}\mathsf{g}_{k}\approx\frac{\zeta_{\mathsf{oc}}\Psi_{k}}{4\pi}\int_{-\frac{M_{z}\epsilon_{k}}{2}}^{\frac{M_{z}\epsilon_{k}}{2}}{\int_{-\frac{M_{x}\epsilon_{k}}{2}}^{\frac{M_{x}\epsilon_{k}}{2}}}f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z,\end{split}$ (A47)
which can be calculated by utilizing [25, Eqs. (2.264.5) & (2.284)]. Furthermore, the expression of $\rho_{\mathsf{s}}$ follows from its definition, which completes the entire proof.

### v2

@@
 The results are further extended to a multiuser case by characterizing the capacity limits of a two-user channel and proposing the associated capacity-achieving decoding and encoding schemes.
-1) For the uplink case, the sum-rate capacity and capacity region, as well as the capacity-achieving detectors, are derived.
-2) For the downlink case, the uplink-downlink duality is established by deriving the uplink-to-downlink and downlink-to-uplink transformations under the same power constraint, based on which the optimal power allocation policy and the achieved sum-rate capacity and capacity region are characterized.
+In the uplink case, the capacity-achieving detectors and sum-rate capacity are derived, and the capacity region is characterized.
+In the downlink case, the uplink-downlink duality is established by deriving the uplink-to-downlink and downlink-to-uplink transformations under the same power constraint, based on which the optimal source current distributions and the achieved sum-rate capacity and capacity region are characterized.
+For comparison, the uplink and downlink sum-rates achieved by the linear zero-forcing scheme are also analyzed.
 To gain further insights, several case studies are presented by specializing the derived results into various array structures, including the planar CAPA, linear CAPA, and planar spatially discrete array (SPDA).
-Numerical results are provided to reveal that: i) the channel capacity achieved by CAPAs converges towards a finite upper bound as the aperture size increases; and ii) CAPAs offer significant capacity gains over the conventional SPDAs.
+Numerical results are provided to reveal that the channel capacity achieved by CAPAs converges towards a finite upper bound as the aperture size increases; and CAPAs offer superior capacity over the conventional SPDAs.
 Index Terms:
-Channel capacity, continuous aperture array (CAPA), performance analysis.
+Channel capacity, continuous-aperture array (CAPA), downlink, performance analysis, uplink.
 I Introduction
-Multiple antenna technology serves as a cornerstone in the advancement of modern wireless communication systems.
-At its core lies the principle of leveraging an increased number of antenna elements to enhance spatial degrees of freedom (DoFs) and augment channel capacity [2].
-Traditionally, multiple antenna prototypes follow a spatially discrete topological structure, where each antenna is depicted as a distinct point [3].
-Motivated by the advantages of incorporating a greater number of antennas, the concept of densely packed antennas has garnered considerable interest in the communication field.
-Within a fixed array aperture, reducing the distance between elements enables the addition of more antennas, thus increasing spatial DoFs.
-This development has led to the emergence of holographic arrays [4, 5].
-In such arrays, antenna elements are densely packed in a configuration that is spaced less than half a wavelength apart, resulting in enhanced spectral efficiency [6].
-Building on the advancements discussed, this paper introduces a novel concept in antenna design—continuous aperture array (CAPA).
-CAPA represents a paradigm shift from the conventional spatially discrete array (SPDA), evolving towards an array that forms a spatially continuous electromagnetic (EM) aperture.
-Essentially, a CAPA can be envisioned as an SPDA comprising an infinite number of antennas with infinitesimal spacing [3, 7].
-The concept of a CAPA marks a significant advancement towards maximizing the spatial DoFs available in antenna array configurations, thereby providing substantial gains in system performance.
-Transitioning from discrete to continuous aperture arrays enhances signal resolution and provides more flexible control over EM wave propagation.
-Through this innovative approach, CAPA paves the way for next-generation antenna technologies, promising to revolutionize wireless communication infrastructures with its advanced spatial processing capabilities [8].
-However, unlike traditional arrays that are modeled using spatially discrete approaches, the modeling of CAPA-based wireless transmission introduces a fundamentally different method grounded in EM field theory.
-Specifically, while the channels of conventional SPDAs are typically represented through discrete matrices, the spatial response of a CAPA necessitates representation through a continuous operator in Hilbert space [9, 10].
-This shift from discrete to continuous modeling is not merely a technical detail but a transformative change in how wireless transmission systems are understood and designed [11].
-The use of continuous operators to model the spatial responses of CAPAs signifies a departure from the traditional matrix-based methods suited to SPDAs.
-This change is crucial because it implies that the entire framework for analyzing and designing transmissions in traditional SPDAs becomes fundamentally inapplicable to CAPAs.
-Consequently, new conceptual and mathematical frameworks tailored specifically to handle the continuous nature of EM field interactions within CAPAs are necessary [9, 10].
-This development is essential for fully exploiting the potential of CAPAs in enhancing system performance.
-Currently, research on CAPA remains in a nascent stage and is predominantly focused on single-user scenarios [11, 12, 13, 14, 15].
-While these foundational studies have laid the groundwork, exploration into the performance of multiuser CAPA systems is considerably less developed.
-Their performance, such as channel capacity, has been seldom analyzed, with only [16] presenting some initial results.
-Specifically, the authors of [16] proposed a sophisticated channel model that utilizes a Green’s function method and leverages a continuous electric current distribution.
-Their research also extends to the performance analysis of CAPA-based reconfigurable intelligent surfaces (RIS) and simultaneously transmitting and reflecting RISs, which provides new insights into the capabilities of CAPAs in more dynamic settings.
-Despite these advancements, a significant gap remains in our understanding of the full potential of CAPAs, particularly regarding their fundamental performance limits.
-More importantly, an analytically tractable framework for CAPA-based communications is lacking.
-This framework is essential for advancing both the theoretical understanding and practical deployment of these systems.
-It would facilitate a deeper comprehension of CAPAs’ capabilities and enable the design of more effective and efficient communication strategies tailored to their unique properties, especially in multiuser scenarios where complexity and demand for system efficiency are significantly higher.
-Motivated by these knowledge gaps, this paper proposes a novel analytical framework to address the continuous nature of EM fields within CAPAs, providing a robust foundation for modeling and understanding CAPA-based communications.
-Using this framework, the channel capacity for both downlink and uplink CAPA-based communications is characterized.
+Multiple-antenna technology is considered one of the fundamental building blocks of modern wireless communication systems.
+At its core lies the principle of utilizing an increased number of antenna elements and expanded array apertures to enhance spatial degrees of freedom (DoFs) and array gains, ultimately improving channel capacity [2, 3].
+In this context, technologies such as multiple-input multiple-output (MIMO) and massive MIMO have played a crucial role in advancing cellular networks from the third generation (3G) to the fourth (4G), fifth (5G), and beyond [4, 5, 6].
+In general, the number of spatial DoFs and array gains scale with the number of antenna elements and the array aperture size.
+Motivated by this principle, various advanced MIMO variants and novel array architectures have been proposed.
+Examples include holographic MIMO [7], gigantic MIMO [8], extremely large aperture arrays [9], reconfigurable intelligent surfaces [10], dynamic metasurface antennas [11], and fluid [12], movable [13], or pinching [14] antennas.
+Extensive theoretical analyses and experimental studies have demonstrated the substantial performance gains achievable through these emerging multiple-antenna technologies.
+Despite their diverse architectures, these novel multiple-antenna systems share a common evolutionary trend: larger aperture sizes, denser antenna configurations, higher operating frequencies, and more flexible structures.
+This progression paves the way for the development of an (approximately) continuous electromagnetic (EM) aperture [15, 16, 17], commonly referred to as a continuous-aperture array (CAPA) [18, 19, 20], which is to be focused on in this paper.
+A CAPA functions as a single, large-aperture antenna with a continuous current distribution, which comprises a (virtually) infinite number of radiating elements integrated with electronic circuits and driven by a limited number of radio-frequency (RF) chains.
+Unlike conventional spatially discrete arrays (SPDAs), a CAPA fully exploits the entire aperture surface and allows for precise control over the current distribution [20].
+This capability significantly enhances spatial DoFs and array gains, as demonstrated in various research endeavors; see Section I-A2 for more relevant references.
+However, the continuous nature of EM field interactions in CAPAs necessitates a fundamental shift in system modeling.
+Instead of the matrix-based models used for SPDAs, CAPA-based systems require an EM-theoretic approach, leading to a continuous Hilbert–Schmidt operator-based integral framework [21, 22, 23, 24, 25, 26, 27].
+This transition from discrete to continuous modeling is not just a mathematical refinement but a paradigm shift in the design and analysis of wireless transmission systems.
+To fully unlock the potential of CAPAs, a novel conceptual framework tailored to their unique characteristics is essential.
+I-A Prior Works
+I-A1 Hardware Implementation of CAPAs
+The development of semi-CAPA and CAPA architectures has a long history in antenna design research, dating back to pioneering contributions in the 1960s [28, 29].
+With recent advancements in materials science and array fabrication, researchers have successfully developed CAPA prototypes based on designs such as metasurface-based leaky-wave antennas [30, 31], optically driven tightly coupled arrays [32], and interdigital transducer-based grating antennas [33]; see [34, 20] for further details.
+Some of these prototypes have even reached commercialization, with early test reports demonstrating their significant potential for enhancing coverage and throughput in practical wireless networks [35, 36].
+It is worth noting that most CAPA prototypes available today are semi-continuous; however, they are effectively treated as continuous antennas, as emphasized by their inventors [35, 36].
+Achieving a true CAPA would require modulating phase and amplitude at each aperture point in an analog manner.
+One possible realization involves using a metasurface composed of materials such as photographic emulsion, which, when illuminated by a tailored wave, could enable unique phase and amplitude modulation at each point.
+This approach would ultimately generate the desired continuous current distribution when a reference wave propagates over the surface [4, 34].
+Such advancements align with the objectives of leading enterprises, which are striving for an advanced level of spatial continuity in antenna design.
+Besides the aforementioned array architectures, several key challenges must be carefully addressed to ensure the practicality and feasibility of CAPA and semi-CAPA systems.
+Some of the most critical challenges include:
+  • EM Coupling: When antennas are densely or continuously packed, mutual coupling (MC) effects become significant.
+MC occurs where EM waves transmitted by one antenna are absorbed by adjacent antennas, which alters their circuitry and distorting the overall spatial channel [37].
+In a CAPA or an ultra-dense semi-CAPA, EM coupling can lead to severe performance degradation, increased interference, and complex interactions that are difficult to predict and mitigate.
+  • Complex Feed Network: A CAPA or an ultra-dense semi-CAPA requires a highly complex feed network to properly excite the aperture or individual elements.
+Designing such a network for a continuous surface or a large number of densely packed elements introduces substantial complexity, increases insertion losses, and raises overall power consumption [38, 39].
+  • Material Limitations: The materials used to construct and sustain a CAPA or an ultra-dense semi-CAPA must meet stringent physical requirements.
+Challenges such as thermal management, mechanical durability, and environmental susceptibility (e.g., humidity, temperature variations, and material aging) impose additional constraints on long-term reliability and scalability [38, 39].
+Taken together, these challenges indicate that realizing a fully functional CAPA or a densely packed semi-CAPA remains elusive, even with recent advancements in materials and fabrication techniques.
+Addressing these issues requires collaborative efforts from both the communication and antenna communities.
+I-A2 Theoretical Design and Analysis of CAPAs
+Although research on the hardware implementation of an ideal CAPA is still in its infancy, increasing efforts have been devoted to the theoretical design and analysis of CAPA- or semi-CAPA-based communications.
+Many of these studies rely on simplified models, such as neglecting the effects of EM coupling and considering narrowband scenarios.
+One of the most widely studied topics is the number of EM DoFs [6], which has been investigated using various analytical tools, including diffraction theory [40], Nyquist-Shannon sampling theorem [41, 21], and Landau’s theorem [42].
+Beyond DoFs, the array gain achieved by CAPAs has also been analyzed in [43].
+These initial studies collectively highlight the advantages of CAPAs over traditional SPDAs.
+In addition, significant research attention has been given to characterizing the channel capacity of CAPA-based point-to-point communication systems.
+These studies leverage Green’s function-based models to describe the EM channel and apply Hilbert operator theory to derive capacity expressions under various propagation environments [23, 24, 22, 26, 25, 27].
+Recently, this research has been extended to scenarios where a CAPA-based reconfigurable surface assists communications between SPDAs [44].
+Another key research direction focuses on continuous beamforming design for CAPA-based communications.
+One prevalent approach relies on discretization-based methods, such as Fourier series expansion, to approximate continuous signals and simplify the associated optimization problems [17, 19].
+Additionally, calculus of variations has been applied to design low-complexity beamforming strategies [45] and explore optimal beamforming structures [46, 47].
+As mentioned earlier, these research developments have relied on simplified system models that primarily consider the spatially continuous nature of CAPAs while omitting other crucial factors, such as the impact of MC and the challenges posed by wideband scenarios.
+To address these limitations, the authors in [48] proposed an EM-based array manifold that accounts for complex physical interactions, including MC.
+This framework enables the modeling of arbitrary antenna configurations, including CAPAs, with greater physical accuracy.
+Regarding wideband transmission, the work in [49] introduced a small-scale fading model for CAPA-based communications using $k$-space (wavenumber domain) analysis.
+This approach provides a wavenumber (spatial-frequency) representation of how the aperture is excited, which illustrates how energy is distributed across different propagation directions (angles) at each frequency.
+Such insights can be instrumental in tackling spatial-wideband effects in CAPA-based wideband communications.
+Another research direction involves circuit-based modeling.
+A recent example in [50] employs multi-port circuit theory combined with Chu’s antenna model to investigate the impact of MC on the wideband behavior of ultra-dense arrays.
+I-B Motivation and Contributions
+Despite these advantages, the fundamental theoretical limits of CAPA-based communications remain largely unexplored.
+Building on existing research in CAPA-based single-user communications, this paper takes a step forward by investigating the capacity limits of CAPA-based multiuser systems and establishing an analytically tractable framework for multiuser CAPA communications.
+Specifically, we consider a scenario where a CAPA serves multiple users equipped with traditional single-antenna devices.
+For simplicity, we focus on a narrowband model, while extending the analysis to wideband scenarios remains an important avenue for future research.
 The main contributions are summarized as follows.
   • We propose a transmission framework for uplink and downlink CAPA communications using EM field theories, where continuous operators model both signals and spatial responses.
-Based on this framework, we design optimal detectors and source current distributions to achieve the single-user channel capacity for uplink and downlink CAPA communications, and derive closed-form expressions for the corresponding channel capacity.
-  • We then extend our analysis to multiuser CAPA systems by studying a two-user case.
-For uplink CAPA communications, we propose using successive decoding to achieve capacity and derive closed-form expressions for the optimal detectors of each user.
+Based on this framework, we design optimal detectors and source current distributions to achieve the single-user channel capacity and derive closed-form expressions for the corresponding capacity.
+  • We extend our analysis to multiuser CAPA systems by studying a two-user case.
+For uplink communications, we propose using successive decoding to achieve capacity and derive closed-form expressions for the optimal detectors of each user.
 Based on this, we derive closed-form expressions for the per-user rate and the sum-rate capacity, and characterize the capacity region.
-For downlink CAPA communications, we establish the uplink-downlink duality by presenting the uplink-to-downlink and downlink-to-uplink transformations under the same sum power constraint and derive closed-form expressions for the capacity-achieving source current distributions.
+For downlink communications, we establish the uplink-downlink duality by presenting the uplink-to-downlink and downlink-to-uplink transformations under the same sum power constraint and derive closed-form expressions for the capacity-achieving source current distributions.
 Additionally, we characterize the downlink capacity limits by deriving the per-user rate, sum-rate capacity, and capacity region.
-  • To unveil further system insights, we present several case studies exploring the line-of-sight (LoS) channel model.
+For comparison, we also analyze the uplink and downlink sum-rates achieved using the linear zero-forcing (ZF) detector and precoder, respectively.
+  • To unveil further insights, we present several case studies exploring the line-of-sight (LoS) channel model.
 The derived results are specialized to various array structures, including the planar CAPA, linear CAPA, and planar SPDA.
@@
 Our results demonstrate that, compared to CAPA, the capacity of the conventional SPDA is limited by its array occupation ratio.
-Moreover, CAPA offers significant capacity enhancements over conventional SPDA, whose capacity converges to that of CAPA as the array occupation ratio approaches one.
+Moreover, CAPA can theoretically achieve greater capacity than conventional SPDA, whose capacity converges to that of CAPA as the array occupation ratio approaches one.
 The remainder of this paper is organized as follows.
 Section II introduces the transmission framework for CAPA-based communications.
-Sections III and IV delve into the capacity analysis and the capacity-achieving schemes for single-user and multiuser CAPA systems, respectively.
+Sections III and IV delve into the capacity analysis and the capacity-achieving schemes for uplink and downlink transmission, respectively.
 Section V provides case studies for several specific array structures.
@@
 […]
-${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})$ $=\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){x}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}}$ (1a)
-$=s_{{\mathsf{ul}},k}\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){\mathsf{J}}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}},$ (1b)
+${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})$ $=\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){x}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}},$ (1)
 […]
@@
 At the BS, the downlink electrical signal is designed as follows:
-${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum\nolimits_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k}$ (4)
+${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum\nolimits_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k},$ (4)
 […]
 ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})=\int_{{\mathcal{A}}}{\mathsf{H}}(\mathbf{s},{\mathbf{r}})\left(\sum\nolimits_{k^{\prime}=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k^{\prime}}({\mathbf{r}})s_{{\mathsf{dl}},k^{\prime}}\right){\rm{d}}{\mathbf{r}}.$ (5)
-Since the aperture size of each user is generally negligible with respect to both the propagation distance and the BS aperture size, the variations of the electric field and channel response along the receive aperture can be omitted.
-This yields ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s}_{k})\triangleq{\mathsf{E}}_{{\mathsf{dl}},k}$ and ${\mathsf{H}}(\mathbf{s},{\mathbf{r}})\approx{\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})={\mathsf{H}}({\mathbf{r}},\mathbf{s}_{k})=\mathsf{H}_{k}(\mathbf{r})$ for $\mathbf{s}\in{\mathcal{A}}_{k}$.
-Therefore, we obtain
+Since the variations of the electric field and channel response along the receive aperture are negligible, we have ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s}_{k})\triangleq{\mathsf{E}}_{{\mathsf{dl}},k}$ and ${\mathsf{H}}(\mathbf{s},{\mathbf{r}})\approx{\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})$ for $\mathbf{s}\in{\mathcal{A}}_{k}$.
+By further assuming channel reciprocity between the uplink and downlink, we have ${\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})={\mathsf{H}}({\mathbf{r}},\mathbf{s}_{k})=\mathsf{H}_{k}(\mathbf{r})$, which gives
 ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}=\!\int_{{\mathcal{A}}}{\mathsf{H}}_{k}({\mathbf{r}})\left(\sum\nolimits_{k^{\prime}=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k^{\prime}}({\mathbf{r}})s_{{\mathsf{dl}},k^{\prime}}\!\right){\rm{d}}{\mathbf{r}}.$ (6)
@@
 II-B Channel Model
-To facilitate theoretical investigations into fundamental performance limits and asymptotic behaviors, we focus our discussion on LoS channels.
-Under this consideration, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ is modeled as follows [19]:
+To facilitate theoretical investigations into fundamental performance limits, we focus our discussion on LoS channels. In this case, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ is modeled as follows [43]:
 ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})={\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}}){\mathsf{H}}^{\mathsf{pa}}(\mathbf{r},{\mathbf{s}}),$ (8)
 […]
-${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\triangleq\frac{{\rm{j}}k_{0}\eta{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}\left(1+\frac{{\rm{j}}k_{0}^{-1}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}\right)$ (9)
-models the influence of free-space EM propagation, where $\eta=120\pi$ is the impedance of free space, and $k_{0}=\frac{2\pi}{\lambda}$ with $\lambda$ being the wavelength denotes the wavenumber.
-Particularly, ${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})$ comprises three terms: the first term corresponds to the radiating near-field and far-field regions, while the remaining two terms correspond to the reactive near-field region.
-The squared magnitude of the parenthesis in (9) is $1-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}+\frac{k_{0}^{-4}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{4}}$, which equals $0.97$ at distance $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert=\lambda$ [19].
-Hence, when considering practical systems with $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert\gg\lambda$, the last two terms in (9) can be neglected.
-As a result, (9) is approximated as follows:
-${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\approx\frac{{\rm{j}}k_{0}\eta}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}.$ (10)
-III Single-User Case
+${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\triangleq\frac{{\rm{j}}k_{0}\eta{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}$ (9)
+models the influence of free-space EM propagation [43], where $\eta=120\pi\,\Omega$ is the impedance of free space, and $k_{0}=\frac{2\pi}{\lambda}$ with $\lambda$ being the wavelength denotes the wavenumber.
+If we consider a rich-scattering environment, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ can be modeled as a stochastic process [7, 49].
+Additionally, the channel model in (8) neglects the impact of MC.
+If MC is taken into account, we could model ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ as follows:
+${\mathsf{H}}(\mathbf{r},{\mathbf{s}})=\int_{{\mathcal{A}}}\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime}){\mathsf{H}}_{\mathsf{nmc}}(\mathbf{r}^{\prime},{\mathbf{s}}){\rm{d}}\mathbf{r}^{\prime},$ (10)
+which generalizes the commonly used linear MC model [37].
+Here, the operator $\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime})$ captures the MC effects within the CAPA, while ${\mathsf{H}}_{\mathsf{nmc}}(\mathbf{r}^{\prime},{\mathbf{s}})$ represents the spatial channel response without considering MC.
+Modeling $\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime})$ requires circuit-theoretic knowledge [50, 37], which is beyond the scope of this paper.
+A detailed study of MC effects on CAPA-based communications is left for future work, although some initial simulation results are provided in Section VI.
+For the rest of this paper, we adopt the LoS channel model as an illustrative example.
+However, it is important to emphasize that our proposed approaches are not constrained to any specific channel type, nor do they rely on assumptions about MC.
+Since all the derived results are based on continuous operator formulations, they remain applicable in a wide range of channel environments.
+III Uplink Capacity
+This section characterizes the channel capacity for CAPA-based uplink communications.
+We commence with the single-user case, and then proceed to explore the two-user case, while extensions to scenarios with an arbitrary number of users will be addressed in future work.
+III-A Single-User Case
 […]
-III-A Uplink Capacity
-We commence with the uplink scenario. In the single-user case, (3b) simplifies to
-${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})=\mathsf{H}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}}}A_{\mathsf{u}}s_{{\mathsf{ul}}}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}).$ (11)
-We then design a detector ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to recover $s_{{\mathsf{ul}}}$ from ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$.
-Specifically, applying this detector to ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ yields
 $\begin{split}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}&=s_{{\mathsf{ul}}}{\mathsf{J}}_{{\mathsf{ul}}}A_{\mathsf{u}}\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\\ &+\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}{\mathbf{r}}.\end{split}$ (11)
@@
 Please refer to Appendix -A for more details. ∎
-The signal-to-noise ratio (SNR) for decoding $s_{{\mathsf{ul}}}$ is thus given by
+The signal-to-noise ratio (SNR) for decoding $s_{{\mathsf{ul}}}$ reads
 $\gamma_{\mathsf{ul}}=\frac{A_{\mathsf{u}}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}}}\rvert^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma^{2}\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}.$ (12)
-The subsequent task is to design ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$ to maximize $\gamma_{\mathsf{ul}}$.
-As observed from (13), $\gamma_{\mathsf{ul}}$ is independent of the norm of ${\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})$, i.e., $\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}$.
-Thus, the optimal detector satisfies
+It can be observed from (12) that the optimal detector satisfies
 ${\mathsf{V}}_{\mathsf{ul}}^{\star}(\mathbf{r})$ $\propto\operatorname*{argmax}_{\int_{{\mathcal{A}}}\lvert{\mathsf{V}}_{\mathsf{ul}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}=1}{\left\lvert\int_{{\mathcal{A}}}{\mathsf{V}}_{\mathsf{ul}}^{*}(\mathbf{r}){\mathsf{H}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\right\rvert^{2}}=\frac{{\mathsf{H}}({\mathbf{r}})}{{\mathsf{h}}_{\mathsf{u}}^{{1}/{2}}},$ (13)
@@
 Remark 1.
-The results in (14) suggest that the capacity-achieving detector for uplink single-user CAPA communications aligns with the spatial channel response ${\mathsf{H}}(\mathbf{r})$, which can be regarded as the continuous version of maximal-ratio combining (MRC).
-Substituting (14) back into (13) gives
+The results in (13) suggest that the capacity-achieving detector for uplink single-user CAPA communications aligns with the spatial channel response ${\mathsf{H}}(\mathbf{r})$, which is the continuous version of maximal-ratio combining (MRC).
+Substituting (13) into (12) gives
 $\gamma_{\mathsf{ul}}=\frac{A_{\mathsf{u}}^{2}\lvert{\mathsf{J}}_{{\mathsf{ul}}}\rvert^{2}}{\sigma^{2}}\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}=\overline{\gamma}_{\mathsf{ul}}\int_{{\mathcal{A}}}\lvert{\mathsf{G}}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}},$ (14)
@@
 ${\mathsf{C}}_{\mathsf{ul}}=\log_{2}(1+\gamma_{\mathsf{ul}})=\log_{2}(1+{\overline{\gamma}_{\mathsf{ul}}}{\mathsf{g}}_{\mathsf{u}}).$ (15)
-III-B Downlink Capacity
-Turning to the downlink case, for $K=1$, (7b) simplifies to
-${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}+{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}),$ (17)
-where ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{\mathsf{u}}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$.
-Applying the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ to ${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})$ yields
-$\begin{split}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}&=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\\ &+\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}.\end{split}$ (18)
-Similar to Lemma 1, we have $\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}\sim{\mathcal{CN}}(0,{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}})$.
-As a result, the SNR for decoding $s_{{\mathsf{dl}}}$ is given by
-$\gamma_{\mathsf{dl}}=\frac{\lvert\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\rvert^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}}.$ (19)
-By noting that ${A_{\mathsf{u}}}$ is negligible with respect to the propagation distance, we have
-$\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\approx{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})A_{\mathsf{u}},\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}\approx\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}},$
-which simplifies (19) as follows:
-$\gamma_{\mathsf{dl}}$ $\approx\frac{\lvert{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}}$ (20a)
-$=\frac{A_{\mathsf{u}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}}.$ (20b)
-The results in (20b) suggest that the downlink SNR is independent of the detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ under our settings.
-Thus, any detector can be used to recover $s_{{\mathsf{dl}}}$.
-Additionally, from (20b), it is observed that the SNR is maximized when the current distribution satisfies ${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\propto{\mathsf{H}}^{*}({\mathbf{r}})$, i.e.,
-${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})=\frac{{\mathsf{H}}^{*}({\mathbf{r}})}{{\mathsf{h}}_{\mathsf{u}}^{1/2}}\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}},$ (21)
-where $\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}$ is proportional to the transmit power.
-Remark 2.
-The results in (21) suggest that the capacity-achieving source current for downlink single-user CAPA communications aligns with ${\mathsf{H}}({\mathbf{r}})$, which can be regarded as the continuous version of maximal-ratio transmission (MRT).
-Inserting (21) into (20b) gives
-$\gamma_{\mathsf{dl}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}{{\sigma}_{\mathsf{u}}^{2}}\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}={\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}},$ (22)
-where ${\overline{\gamma}_{\mathsf{dl}}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}k_{0}^{2}\eta^{2}}{4\pi{\sigma}_{\mathsf{u}}^{2}}$ can be interpreted as the downlink transmit SNR.
-Based on (22), the single-user downlink capacity is given by
-${\mathsf{C}}_{\mathsf{dl}}=\log_{2}(1+\gamma_{\mathsf{dl}})=\log_{2}(1+{\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}}).$ (23)
-IV Multiuser Case
-This section extends the single-user scenario to the multiuser one.
-We characterize the channel capacity for both uplink and downlink communications.
-For brevity, our analysis focuses on the two-user case, while extensions to scenarios with an arbitrary number of users will be addressed in future work.
-IV-A Uplink Capacity
-We first analyze the uplink scenario.
-IV-A1 Optimal Uplink Detection
+III-B Two-User Case
+We now extend the single-user scenario to the two-user one.
+III-B1 Optimal Uplink Detection
 When $K=2$, the observed electric field in (3b) is given by
@@
 ${\mathsf{Z}}(\mathbf{r})\triangleq\mathsf{H}_{1}(\mathbf{r}){\mathsf{J}}_{{\mathsf{ul}},1}A_{{\mathsf{u}},1}s_{{\mathsf{ul}},1}+{\mathsf{N}}_{\mathsf{ul}}(\mathbf{r}),$ (17)
-which is the same as that given in (11).
 Following the derivation steps to obtain (13), the BS can use the detector ${\mathsf{V}}_{{\mathsf{ul}},1}(\mathbf{r})=\frac{{\mathsf{H}_{1}}({\mathbf{r}})}{\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{H}}_{1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}}$ to recover $s_{{\mathsf{ul}},1}$ from (17).
@@
   1. i
-    Whitening: The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Z}}(\mathbf{r}){\rm{d}}\mathbf{r}$ is proportional to the Dirac delta function, i.e.,
+    Whitening: The autocorrelation function of ${\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Z}}(\mathbf{r}){\rm{d}}\mathbf{r}$ aligns with the delta function, i.e.,
     ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}\propto\delta(\mathbf{r}^{\prime}-\mathbf{r});$ (20)
@@
 Please refer to Appendix -B for more details. ∎
-Next, we exploit ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to transform ${\mathsf{Y}}_{\mathsf{B}}(\mathbf{r})$ as follows:
+Next, we exploit ${\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r})$ to transform ${\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r})$ as follows:
 $\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{Y}}_{\mathsf{ul}}(\mathbf{r}){\rm{d}}\mathbf{r}=\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime}){\mathsf{J}}_{{\mathsf{ul}},2}A_{{\mathsf{u}},2}s_{{\mathsf{ul}},2}+{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}^{\prime}),$ (23)
 where $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})\triangleq\int_{{\mathcal{A}}}{\mathsf{W}}_{{\mathsf{Z}}}(\mathbf{r}^{\prime},\mathbf{r}){\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}\mathbf{r}$.
-By noting that ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$, the model presented in (31) is similar to that in (11).
+By noting that ${\mathbbmss{E}}\{{\mathsf{Z}}_{\mathsf{w}}(\mathbf{r}){\mathsf{Z}}_{\mathsf{w}}^{*}(\mathbf{r}^{\prime})\}={\sigma}^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})$, the model presented in (23) is similar to that in the single-user case.
 Motivated by this, we design an MRC-based detector $\mathsf{V}_{\mathsf{ul},2}(\mathbf{r}^{\prime})$ by treating $\overline{\mathsf{H}}_{2}(\mathbf{r}^{\prime})$ as the equivalent channel response, which yields
@@
 TABLE I: SIC decoding for uplink CAPA communications.
-IV-A2 Capacity Characterization
+III-B2 Capacity Characterization
 After obtaining the decoding SNR of each $s_{{\mathsf{ul}},k}$ for $k=1,2$, the uplink achievable rate of user $k$ can be calculated as ${\mathsf{R}}_{\mathsf{ul},k}^{2\rightarrow 1}=\log_{2}(1+\gamma_{\mathsf{ul},k})$, i.e.,
@@
 This region can be achieved through SIC decoding along with time sharing, which forms a pentagon.
-IV-B Downlink Capacity
-In this part, we focus on the downlink scenario.
+III-B3 Sum-Rate Achieved by Zero-Forcing Detector
+To attain the uplink sum-rate capacity in (33), a non-linear detector based on SIC is necessary.
+However, in practical systems, the linear ZF detector is frequently favored to lower decoding complexity.
+In the subsequent analysis, we aim to evaluate the sum-rate performance of the ZF detector and contrast it with the sum-rate capacity achieved by the non-linear detector.
+Let $\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})$ denote the ZF detector for user $k$ with $k\in\{1,2\}$.
+In order to nullify the IUI, i.e., $\int_{{\mathcal{A}}}\mathsf{V}_{\mathsf{zf},k}^{*}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r}){\rm{d}}{\mathbf{r}}=0$ for $k^{\prime}\neq k$, and maximize the channel gain of user $k$, $\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})$ is designed as the projection of $\mathsf{H}_{k}(\mathbf{r})$ onto the space orthogonal to $\mathsf{H}_{k^{\prime}}(\mathbf{r})$, which is given by
+$\mathsf{V}_{\mathsf{zf},k}(\mathbf{r})=\mathsf{H}_{k}(\mathbf{r})-\frac{\int_{\mathcal{A}}{\mathsf{H}_{k}}(\mathbf{r})\mathsf{H}_{k^{\prime}}^{*}(\mathbf{r})\mathrm{d}\mathbf{r}}{\int_{\mathcal{A}}{\left|\mathsf{H}_{k^{\prime}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}\mathsf{H}_{k^{\prime}}(\mathbf{r}).$ (36)
+Substituting (36) into $\mathsf{R}_{\mathsf{ul},k}^{\mathsf{zf}}=\log_{2}(1+\gamma_{\mathsf{ul},k}^{\mathsf{zf}})$ with
+and performing some simple manipulations, we can obtain the sum-rate achieved by the ZF detector as follows:
+$\mathsf{R}_{\mathsf{ul}}^{\mathsf{zf}}=\sum\nolimits_{k=1}^{2}{\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{ul},k}\mathsf{g}_{k}\big{(}1-\left|\rho\right|^{2}\big{)}\Big{)}}.$ (37)
+From the results, it can be observed that $\mathsf{R}_{\mathsf{ul}}^{\mathsf{zf}}\leq\mathsf{C}_{\mathsf{ul}}$, and the gap between them is further illustrated in the numerical results presented in Section VI.
+IV Downlink Capacity
+In this section, we characterize the channel capacity for CAPA-based downlink communications.
+IV-A Single-User Case
+For $K=1$, (7b) simplifies to
+${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}+{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}),$ (38)
+where ${\mathbbmss{E}}\{{\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}^{*}({\mathbf{s}}^{\prime})\}={\sigma}_{\mathsf{u}}^{2}\delta(\mathbf{s}-{\mathbf{s}}^{\prime})$.
+Applying a detector ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$ to ${\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s})$ yields
+$\begin{split}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{Y}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}&=s_{{\mathsf{dl}}}\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\\ &+\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}.\end{split}$ (39)
+Similar to Lemma 1, we have $\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\mathsf{N}}_{{\mathsf{dl}}}(\mathbf{s}){\rm{d}}{\mathbf{s}}\sim{\mathcal{CN}}(0,{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}})$.
+As a result, the SNR for decoding $s_{{\mathsf{dl}}}$ is given by
+$\gamma_{\mathsf{dl}}=\frac{\lvert\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\rvert^{2}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}}.$ (40)
+Recalling that the variations of the signal along ${{\mathcal{A}}_{\mathsf{u}}}$ are negligible, we have $\int_{{\mathcal{A}}_{\mathsf{u}}}{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}){\rm{d}}{\mathbf{s}}\approx{\mathsf{V}}_{{\mathsf{dl}}}^{*}(\mathbf{s}_{\mathsf{u}})A_{\mathsf{u}}$ and $\int_{{\mathcal{A}}_{\mathsf{u}}}\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})\rvert^{2}{\rm{d}}{\mathbf{s}}\approx\lvert{\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s}_{\mathsf{u}})\rvert^{2}A_{\mathsf{u}}$, which simplifies (40) as follows:
+$\gamma_{\mathsf{dl}}\approx\frac{A_{\mathsf{u}}\lvert\int_{{\mathcal{A}}}{\mathsf{H}}({\mathbf{r}}){\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}{{\sigma}_{\mathsf{u}}^{2}}.$ (41)
+We note that $\gamma_{\mathsf{dl}}$ does not rely on ${\mathsf{V}}_{{\mathsf{dl}}}(\mathbf{s})$.
+Moreover, it is observed from (41) that the SNR is maximized when the current distribution satisfies ${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\propto{\mathsf{H}}^{*}({\mathbf{r}})$, i.e.,
+${\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})=\frac{{\mathsf{H}}^{*}({\mathbf{r}})}{{\mathsf{h}}_{\mathsf{u}}^{1/2}}\sqrt{\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}},$ (42)
+where $\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}$ is proportional to the transmit power.
+Remark 3.
+The results in (42) suggest that the capacity-achieving source current for downlink single-user CAPA communications aligns with ${\mathsf{H}}({\mathbf{r}})$, which can be regarded as the continuous version of maximal-ratio transmission (MRT).
+Inserting (42) into (41) gives
+$\gamma_{\mathsf{dl}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}}{{\sigma}_{\mathsf{u}}^{2}}\int_{\mathcal{A}}\lvert{\mathsf{H}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}={\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}},$ (43)
+where ${\overline{\gamma}_{\mathsf{dl}}}=\frac{A_{\mathsf{u}}\int_{{\mathcal{A}}}\lvert{\mathsf{J}}_{{\mathsf{dl}}}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}k_{0}^{2}\eta^{2}}{4\pi{\sigma}_{\mathsf{u}}^{2}}$ can be interpreted as the downlink transmit SNR.
+Based on (43), the single-user downlink capacity is given by
+${\mathsf{C}}_{\mathsf{dl}}=\log_{2}(1+\gamma_{\mathsf{dl}})=\log_{2}(1+{\overline{\gamma}_{\mathsf{dl}}}{\mathsf{g}}_{\mathsf{u}}).$ (44)
+IV-B Two-User Case
+In this part, we focus on the two-user scenario.
 IV-B1 Optimal Multiuser Precoding
@@
 where ${\mathsf{v}}_{\varepsilon_{k}}=\int_{\mathcal{A}_{\varepsilon_{k}}}{\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}(\mathbf{s})\rvert^{2}}\mathrm{d}\mathbf{s}$ and $\overline{\mathsf{v}}_{\varepsilon_{k}}=\lvert\int_{\mathcal{A}_{\varepsilon_{k}}}{\mathsf{V}}_{{\mathsf{dl}},\varepsilon_{k}}^{*}(\mathbf{s})\mathrm{d}\mathbf{s}\rvert^{2}$.
-Since each user’s array aperture is negligible with respect to the propagation distance, we have
+Since the variations of the signal along each user’s aperture is negligible, we have
 $\overline{\mathsf{v}}_{\varepsilon_{k}}\approx\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{s}}_{\varepsilon_{k}})\rvert^{2}A_{{\mathsf{u}},\varepsilon_{k}}^{2},{\mathsf{v}}_{\varepsilon_{k}}\approx\lvert\mathsf{V}_{{\mathsf{dl}},\varepsilon_{k}}({\mathbf{s}}_{\varepsilon_{k}})\rvert^{2}A_{{\mathsf{u}},\varepsilon_{k}},$ (46)
@@
 $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}=\max\nolimits_{\sum_{k=1}^{2}{\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})\rvert^{2}}\mathrm{d}\mathbf{r}}\leq{\mathsf{P}}}\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}+\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1},$ (50)
-where $\mathsf{P}$ reflects the transmit power.
+where $\mathsf{P}$ reflects the power budget.
 Following the same derivation approach, the sum-rate achieved by the DPC order $1\rightarrow 2$ can also be derived.
-For Gaussian channels, the maximal DPC sum-rate is independent of the DPC decoding order, and thus $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}$ is essentially the downlink sum-rate capacity [21].
+For downlink Gaussian channels, the maximal DPC sum-rate is independent of the DPC encoding order, and thus $\mathsf{C}_{\mathsf{dl}}^{2\rightarrow 1}$ is essentially the sum-rate capacity [52].
 Problem (50) is non-convex, making the numerical identification of the maximum a nontrivial task.
@@
 By defining $\overline{\gamma}_{\mathsf{du},k}\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}$ as the transmit SNRs of the dual uplink channel for $k=1,2$, the downlink sum-rate capacity is equal to the following dual uplink capacity [51, 52, 53]:
-$\begin{split}{\mathsf{C}}_{\mathsf{du}}=\max_{{\mathsf{P}}_{1},{\mathsf{P}}_{2}\geq 0,{\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq{\mathsf{P}}}&\log_{2}(1+\overline{\gamma}_{\mathsf{du},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{2}\\ &+\overline{\gamma}_{\mathsf{du},1}\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}),\end{split}$ (50)
+$\begin{split}{\mathsf{C}}_{\mathsf{du}}=\max_{{\mathsf{P}}_{1},{\mathsf{P}}_{2}\geq 0,{\mathsf{P}}_{1}+{\mathsf{P}}_{2}\leq{\mathsf{P}}}\log_{2}(1&+\overline{\gamma}_{\mathsf{du},1}\mathsf{g}_{1}+\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{2}\\ &+\overline{\gamma}_{\mathsf{du},1}\overline{\gamma}_{\mathsf{du},2}\mathsf{g}_{1}\mathsf{g}_{2}\overline{\rho}),\end{split}$ (51)
 which is a convex problem and can be solved as follows.
@@
 As shown in Appendix -D and Section IV-B2, the output source currents achieve the same rates as the dual uplink channel using the input power allocation policy and SIC decoding with the specified decoding order.
-Having obtained the capacity-achieving source currents under the DPC order $2\rightarrow 1$, we summarize the entire precoding procedure in Table II for ease of reference.
-The presented DPC procedure can be directly extended to the other order $1\rightarrow 2$ by exchanging the user indices.
+[…]
 1:Encode the data information into the symbols $\{s_{{\mathsf{dl}},k}\}_{k=1}^{2}$ using DPC with the order $2\rightarrow 1$; 2:Determine the dual power allocation policy $\{{\mathsf{P}}_{k}\}_{k=1}^{2}$ according to the specific target quality of service requirements; 3:Generate the source current ${\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})$: ${\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})=\sqrt{{\mathsf{P}}_{1}}\frac{\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})-\frac{{\mathsf{P}}_{2}\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})}{\sqrt{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{1}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}-\frac{{\mathsf{P}}_{2}\lvert\int_{\mathcal{A}}\hat{\mathsf{H}}_{1}^{*}(\mathbf{r})\hat{\mathsf{H}}_{2}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{1+{\mathsf{P}}_{2}\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}}}};$ 4:Generate the source current ${\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})$: ${\mathsf{J}}_{{\mathsf{dl}},2}({\mathbf{r}})=\sqrt{{\mathsf{P}}_{2}}\frac{\hat{\mathsf{H}}_{2}^{*}(\mathbf{r})\sqrt{1+\lvert\int_{\mathcal{A}}\hat{\mathsf{\mathsf{H}}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\rvert^{2}}}{\int_{\mathcal{A}}\lvert\hat{\mathsf{H}}_{2}(\mathbf{r})\rvert^{2}{\rm{d}}{\mathbf{r}}};$ 5:Formulate the transmit signal ${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k}$.
-TABLE II: DPC-based precoding for downlink CAPA communications.
+TABLE II: DPC for downlink CAPA communications.
 IV-B2 Capacity Characterization
-As proved in Appendix -D, given the power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, the achievable rates of user $1$ and user $2$ under the DPC order $2\rightarrow 1$ are given as follows:
+As proved in Appendix -D, given the power allocation policy $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$, the achievable rates of user $1$ and user $2$ under the DPC order $2\rightarrow 1$ are given by
 $\mathsf{R}_{\mathsf{dl},1}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},1}\mathsf{g}_{1}\left(1-\frac{\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}|\rho|^{2}}{1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}}\right)\right),$ (55)
-$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right).$ (55)
-By comparing the above two equations with the results in Corollary 1, the following observations are made.
+$\mathsf{R}_{\mathsf{dl},2}^{2\rightarrow 1}$ $=\log_{2}\left(1+\overline{\gamma}_{\mathsf{dl},2}\mathsf{g}_{2}\right),$ (56)
+respectively, where $\overline{\gamma}_{\mathsf{dl},k}=\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}^{\star}$ for $k\in\{1,2\}$.
+By comparing the above two expressions with the results in Corollary 1, the following observations are made.
 Remark 5.
 The downlink user rates achieved by the DPC order $2\rightarrow 1$ equal those achieved in the dual uplink channel under the SIC decoding order $1\rightarrow 2$.
-The achievable rates under the DPC order $1\rightarrow 2$ can be derived with similar steps, which are omitted for brevity.
-For a thorough study, we derive the downlink-to-uplink transformation as follows.
+The rates achieved by the DPC order $1\rightarrow 2$ can be derived with similar steps, which are omitted for brevity. In the sequel, we derive the downlink-to-uplink transformation.
 Theorem 4.
@@
 $\mathsf{C}_{\mathsf{dl}}$ $=\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{dl},1}\overline{\gamma}_{\mathsf{dl},2}{\mathsf{g}}_{1}{\mathsf{g}}_{2}\overline{\rho}+\sum\nolimits_{k=1}^{2}\overline{\gamma}_{\mathsf{dl},k}{\mathsf{g}}_{k}\Big{)}$ (59a)
-=$=\begin{cases}\log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},1}(\mathsf{P})\mathsf{g}_{1}\right)&\xi\geq{\mathsf{P}}\\ \log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},2}(\mathsf{P})\mathsf{g}_{2}\right)&\xi\leq-{\mathsf{P}}\\ \log_{2}\left(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{1}\varepsilon_{2}\overline{\rho}\right)&\mathrm{else}\\ \end{cases},$ (58b)
-where $\widetilde{\gamma}_{\mathsf{dl},k}(x)\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}x$ for $k=1,2$, $\varepsilon_{1}=\widetilde{\gamma}_{\mathsf{dl},1}(\frac{{\mathsf{P}}-\xi}{2})\mathsf{g}_{1}$, and $\varepsilon_{2}=\widetilde{\gamma}_{\mathsf{dl},2}(\frac{{\mathsf{P}}+\xi}{2})\mathsf{g}_{2}$.
+=$=\begin{cases}\log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},1}(\mathsf{P})\mathsf{g}_{1}\right)&\xi\geq{\mathsf{P}}\\ \log_{2}\left(1+\widetilde{\gamma}_{\mathsf{dl},2}(\mathsf{P})\mathsf{g}_{2}\right)&\xi\leq-{\mathsf{P}}\\ \log_{2}\left(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{1}\varepsilon_{2}\overline{\rho}\right)&\mathsf{else}\\ \end{cases},$ (59b)
+where $\widetilde{\gamma}_{\mathsf{dl},k}(x)\triangleq\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}x$ for $k\in\{1,2\}$, $\varepsilon_{1}=\widetilde{\gamma}_{\mathsf{dl},1}(\frac{{\mathsf{P}}-\xi}{2})\mathsf{g}_{1}$, and $\varepsilon_{2}=\widetilde{\gamma}_{\mathsf{dl},2}(\frac{{\mathsf{P}}+\xi}{2})\mathsf{g}_{2}$.
 Proof:
@@
 This expression is applicable to an arbitrary aperture, regardless of its location, shape, and size.
-Furthermore, according to the uplink-downlink duality, given a power allocation scheme $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ with ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, we can obtain the capacity region of a dual uplink channel.
+According to the uplink-downlink duality, given a power allocation scheme $({\mathsf{P}}_{1},{\mathsf{P}}_{2})$ with ${\mathsf{P}}_{1}+{\mathsf{P}}_{2}={\mathsf{P}}$, we can obtain the capacity region of a dual uplink channel.
 The downlink capacity region is the convex hull of the union of all these dual uplink capacity regions, which is illustrated in Section VI.
+IV-B3 Sum-Rate Achieved by Zero-Forcing Precoder
+Although DPC is a potent capacity-achieving scheme, its practical implementation proves challenging.
+Consequently, non-DPC linear ZF precoding schemes hold practical significance due to the lower complexity.
+In the following analysis, we aim to evaluate the sum-rate performance achieved by the ZF precoder.
+Let $\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})$ denote the ZF source current for user $k$ with $k\in\{1,2\}$.
+In order to nullify the IUI, i.e., $\int_{{\mathcal{A}}}\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r}){\rm{d}}{\mathbf{r}}=0$ for $k^{\prime}\neq k$, and maximize the channel gain of user $k$, $\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})$ is designed as the projection of $\mathsf{H}^{*}_{k}(\mathbf{r})$ onto the space orthogonal to $\mathsf{H}^{*}_{k^{\prime}}(\mathbf{r})$, which is given by
+$\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})={\mathsf{H}_{k}^{*}(\mathbf{r})}-\frac{\int_{\mathcal{A}}{\mathsf{H}_{k}^{*}}(\mathbf{r})\mathsf{H}_{k^{\prime}}(\mathbf{r})\mathrm{d}\mathbf{r}}{\int_{\mathcal{A}}{\left|\mathsf{H}_{k^{\prime}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}\mathsf{H}_{k^{\prime}}^{*}(\mathbf{r}).$ (60)
+By inserting (60) into the following per-user rate expression
+$\mathsf{R}_{\mathsf{dl},k}^{\mathsf{zf}}=\log_{2}\left(1+\frac{|\int_{\mathcal{A}}{}\mathsf{H}_{k}(\mathbf{r})\mathsf{J}_{\mathsf{zf},k}(\mathbf{r})\mathrm{d}\mathbf{r}|^{2}}{\frac{\sigma_{k}^{2}}{A_{\mathsf{u},k}}+|\int_{\mathcal{A}}{}\mathsf{H}_{k}(\mathbf{r})\mathsf{J}_{\mathsf{zf},k^{\prime}}(\mathbf{r})\mathrm{d}\mathbf{r}|^{2}}\right),$ (61)
+the sum-rate achieved by the ZF precoding is given as follows:
+$\mathsf{R}_{\mathsf{dl}}^{\mathsf{zf}}=\sum\nolimits_{k=1}^{2}{\log_{2}\Big{(}1+\overline{\gamma}_{\mathsf{dl},k}^{\mathsf{zf}}\mathsf{g}_{k}\big{(}1-\left|\rho\right|^{2}\big{)}\Big{)}},$ (62)
+where $\overline{\gamma}_{\mathsf{dl},k}^{\mathsf{zf}}=\frac{A_{{\mathsf{u}},k}k_{0}^{2}\eta^{2}}{4\pi\sigma_{k}^{2}}{\mathsf{P}}_{k}^{\mathsf{zf}}$ with ${\mathsf{P}}_{k}^{\mathsf{zf}}=\int_{\mathcal{A}}{\lvert{\mathsf{J}}_{{\mathsf{zf}},k}(\mathbf{r})\rvert^{2}}\mathrm{d}\mathbf{r}$ being the optimally allocated power that can be obtained by the water-filling algorithm [2, 3].
+V Case Studies
+As discussed in Remark 2 and Remark 8, the sum-rate capacity expressions derived for both uplink and downlink CAPA communications are applicable to apertures of arbitrary location, shape, and size, among other factors.
+In this section, we specialize ${\mathcal{A}}$ to several specific structures to unveil more system insights under the two-user scenario.
 (a) Planar CAPA.
 (b) Linear CAPA.
-Figure 2: Geometries of CAPAs.
-V Case Studies
-As discussed in Remark 3 and Remark 8, the sum-rate capacity expressions derived for both uplink and downlink CAPA communications are applicable to apertures of arbitrary location, shape, and size, among other factors.
-In the following sections, we specialize ${\mathcal{A}}$ to several specific cases to unveil more system insights.
+Figure 2: Illustration of the geometries of CAPAs.
 V-A Planar CAPAs
@@
 The edges of $\mathcal{A}$ are parallel to the axes, with physical dimensions $L_{x}$ and $L_{z}$ along the $x$- and $z$-axes, respectively.
-In this case, we have ${\mathcal{A}}=\{[x,0,z]^{\mathsf{T}}|x,z\in[-\frac{L_{x}}{2},\frac{L_{x}}{2}]\}$.
+In this case, we have $\mathcal{A}=\{[x,0,z]^{\mathsf{T}}|x\in[-\frac{L_{x}}{2},\frac{L_{x}}{2}],z\in[-\frac{L_{z}}{2},\frac{L_{z}}{2}]\}$.
 For each user $k\in\{1,2\}$, let $r_{k}$ represent the distance from the center of $\mathcal{A}$ to the center of $\mathcal{A}_{\mathsf{k}}$, $\phi_{k}\in[0,\pi]$ and $\theta_{k}\in[0,\pi]$ denote the associated azimuth and elevation angles, respectively.
@@
 $\begin{split}\mathsf{G}_{k}(\mathbf{r})&=\frac{\sqrt{r_{k}\Psi_{k}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{1}{2}}}}{\sqrt{4\pi}(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{\frac{3}{4}}}\\ &\triangleq\mathsf{Q}_{k}(x,z).\end{split}$ (63)
-To characterize the capacity, we derive closed-form expressions of the channel gain and correlation factor as follows.
+To calculate the capacity, we derive closed-form expressions of the channel gain and correlation factor as follows.
 Lemma 4.
@@
 The associated channel correlation factor is approximated by
-$\begin{split}\rho=&\frac{\pi^{2}A}{4n^{2}\sqrt{\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}{\sqrt{\left(1-\psi_{j}^{2}\right)\left(1-\psi_{j^{\prime}}^{2}\right)}}\\ &\times\mathsf{Q}_{1}^{*}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\triangleq\rho_{\mathsf{p}},\end{split}$ (61)
+$\begin{split}\rho=&\frac{\pi^{2}A}{4n^{2}\sqrt{\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}{\sqrt{\left(1-\psi_{j}^{2}\right)\big{(}1-\psi_{j^{\prime}}^{2}\big{)}}}\\ &\times\mathsf{Q}_{1}^{*}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\triangleq\rho_{\mathsf{p}},\end{split}$ (65)
 where $n$ is a complexity-vs-accuracy tradeoff parameter, and $\psi_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$ for $j\in\{1,\ldots,n\}$.
@@
 Please refer to Appendix -F for more details. ∎
+Note that the expression in (65) is derived using the Chebyshev–Gauss quadrature rule, as per Appendix -F.
 Subsequently, the closed-form expressions of sum-rate capacity can be readily obtained.
@@
 In this case, we have $\lim_{L_{x},L_{z}\rightarrow\infty}{\mathsf{g}}_{k}^{\mathsf{p}}=\frac{1}{4\pi}\frac{4\pi}{2}=\frac{1}{2}$.
-Regarding $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}$, while it is computationally intractable, based on the numerical results presented in [23, 24], we observe that $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}\ll 1$.
+Regarding $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert$, while it is computationally intractable, based on the numerical results presented in [54, 55], we observe that $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert\ll 1$.
 Taken together, we can obtain the asymptotic sum-rate capacity as follows:
@@
 where $i\in\{\mathsf{ul},\mathsf{dl}\}$.
-The results in (63) suggest that when the BS is equipped with an infinitely large planar CAPA, half of the transmission power of each user can be captured by the receiver, and the inter-user interference vanishes.
+The results in (67) suggest that when the BS is equipped with an infinitely large planar CAPA, half of the transmission power of each user can be captured by the receiver, and the IUI vanishes.
 Remark 9.
 The above observation makes intuitive sense because an infinitely large planar CAPA can receive at most half of the power transmitted by an isotropic source, while the other half of the power will never reach the surface.
+Additionally, since $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert\ll 1$, we have
+$\lim_{L_{x},L_{z}\rightarrow\infty}\mathsf{R}_{i}^{\mathsf{zf}}\approx\lim_{L_{x},L_{z}\rightarrow\infty}\mathsf{C}_{i},$ (68)
+for $i\in\{\mathsf{ul},\mathsf{dl}\}$. This suggest that when the size of the CAPA is sufficiently large, the sum-rates achieved by the ZF schemes can approach the sum-rate capacities.
 V-B Linear CAPAs
@@
 where $i\in\{\mathsf{ul},\mathsf{dl}\}$.
-By comparing (63) with (66), we observe that the asymptotic sum-rate capacity achieved by the linear CAPA is influenced not only by the transmit SNR but also by the locations of the users, which is in contrast to the planar CAPA.
-Figure 3: Planar SPDA.
+By comparing (67) with (71), it is found the asymptotic sum-rate capacity achieved by the linear CAPA is influenced not only by the SNR but also by the locations of the users, which is in contrast to the planar CAPA.
+Figure 3: Illustration of a planar SPDA.
 V-C Planar SPDAs
-Finally, we consider a case where the above planar CAPA is partitioned into $M=M_{z}M_{x}$ spatially discrete elements, where $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$ denote the number of antenna elements along the $x$- and $z$-axes, as depicted in Fig.
+Finally, we consider a planar SPDA with the same size as the planar CAPA, i.e., $L_{x}\times L_{z}$, comprising $M=M_{z}M_{x}$ spatially discrete elements, where $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$ denote the number of antenna elements along the $x$- and $z$-axes, as depicted in Fig.
 3.
@@
 The capacity achieved by an SPDA increases to that achieved by a CAPA when $\zeta_{\mathsf{oc}}=1$. This makes intuitive sense as an SPDA turns into a CAPA when the array occupation ratio equals one.
+We also note that the above conclusion is derived without considering the impact of EM coupling.
+When MC is taken into account, the relationship between the sum-rates achieved by CAPAs and SPDAs may depend on the size of each discrete element as well as the operating frequency [50].
+(a) $\left|\rho\right|^{2}$ vs. $n$ for small aperture sizes.
+(b) $\left|\rho\right|^{2}$ vs. $n$ with $A=10^{6}$ m2.
+Figure 4: Convergence of Chebyshev-Gauss quadrature.
+Figure 5: Uplink transmission rates of CAPA.
 VI Numerical Results
-In this section, computer simulations are performed to demonstrate the performance of CAPA communications and verify the accuracy of the developed analytical results.
+In this section, computer simulations are performed to demonstrate the performance of CAPA-based communications and verify the accuracy of the developed analytical results.
 For clarity, the simulations employ the planar arrays specified in Section V.
-Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.0107$ m, $d=\frac{\lambda}{2}$, $A_{\mathsf{u},1}=A_{\mathsf{u},2}=\frac{\lambda^{2}}{4\pi}$, $L_{x}=L_{z}$, $M_{x}=M_{z}$, $\overline{\gamma}_{\mathsf{ul},1}=30$ dB, $\overline{\gamma}_{\mathsf{ul},2}=40$ dB, $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}=50$ dB with $\sigma_{1}^{2}=\sigma_{2}^{2}$, $\theta_{1}=\theta_{2}=\frac{\pi}{6}$, $\phi_{1}=\phi_{2}=\frac{\pi}{3}$, $r_{1}=10$ m, and $r_{2}=20$ m.
-All the results are calculated by setting the complexity-vs-accuracy tradeoff parameter $n$ as $n=200$.
-Figure 4: Uplink transmission rates achieved by CAPA.
-(a) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
-(b) Sum-rate capacity vs. $A$.
+Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.125$ m, $A_{\mathsf{u},1}=A_{\mathsf{u},2}=A_{\mathsf{s}}=\frac{\lambda^{2}}{4\pi}$, $L_{x}=L_{z}=0.5$ m2, $M_{x}=M_{z}$, $\overline{\gamma}_{\mathsf{ul},1}=30$ dB, $\overline{\gamma}_{\mathsf{ul},2}=40$ dB, $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}=50$ dB with $\sigma_{1}^{2}=\sigma_{2}^{2}$, $\theta_{1}=\theta_{2}=\frac{\pi}{6}$, $\phi_{1}=\phi_{2}=\frac{\pi}{3}$, $r_{1}=10$ m, and $r_{2}=20$ m.
+To select an appropriate value for the complexity-vs-accuracy tradeoff parameter $n$ in the Chebyshev-Gauss quadrature given in (65), we first analyze the convergence behavior of $\left|\rho\right|^{2}$ with respect to $n$ for different aperture sizes, as illustrated in Figure 4.
+The results indicate that for all aperture sizes depicted in Figure 4(a), $\left|\rho\right|^{2}$ stabilizes once $n$ exceeds $20$.
+This means that using $20$ terms for summation in (65) can precisely approximate the integral $\int_{{\mathcal{A}}}{\mathsf{G}}_{1}^{*}({\mathbf{r}}){\mathsf{G}}_{2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}$ contained in the channel correlation factor $\left|\rho\right|^{2}=\frac{\left|\int_{{\mathcal{A}}}{\mathsf{G}}_{1}^{*}({\mathbf{r}}){\mathsf{G}}_{2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right|^{2}}{{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}}$.
+Additionally, we observe that larger aperture sizes require higher values of $n$ for the Chebyshev-Gauss quadrature to converge.
+To ensure convergence for extremely large array sizes, we conduct further simulations by examining an aperture size of $10^{6}$ m2 in Figure 4(b).
+The results reveal that convergence is effectively achieved when $n=1000$.
+Based on these observations and to maintain a balance between computational complexity and accuracy, we set $n=1000$ for Figure 6(a) and 9(a) to illustrate asymptotic rate behavior for extremely larger arrays, while using $n=20$ for the other results.
+(a) Sum-rate capacity vs. $A$.
+(b) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
 Figure 6: Uplink sum-rate capacity.
 VI-A Uplink CAPA Communications
-Figure 4 illustrates the uplink per-user rate and sum-rate capacity achieved by the CAPA in terms of the BS aperture size under the SIC decoding orders $1\rightarrow 2$ and $2\rightarrow 1$.
+Figure 5 illustrates the uplink per-user rate, sum-rate capacity, and the sum-rate achieved by the ZF detector in terms of the BS aperture size.
 It can be observed from this graph that as the aperture size $A$ grows, all the presented rates increase.
 This increase is attributed to the larger aperture size enhancing channel gain and reducing channel correlation, which in turn enhances the dedicated signal and diminishes the IUI, as discussed in (67).
-To compare the performance gap between CAPA and the conventional SPDA in terms of sum-rate capacity, we plot the sum-rate capacity for SPDAs as a function of the array occupation ratio $\zeta_{\mathsf{oc}}$ in Figure 5(a).
-As shown in the graph, the sum-rate capacity for an SPDA gradually converges with that of a CAPA as the array occupation ratio increases towards $1$.
-This observation validates the statements in Remark 10.
-Additionally, the results in Figure 5(a) indicate that increasing the aperture size can boost the channel capacity.
-To gain further insights into the capacity limit as the BS aperture size $A$ approaches infinity, we plot the sum-rate capacity against the aperture size in Figure 5(b).
+Furthermore, it is important to note that the sum-rate achieved by the linear ZF detector is lower than the sum-rate capacity achieved by the proposed optimal SIC-based detector.
+To gain further insights into the capacity limit as the BS aperture size $A$ approaches infinity, we plot the sum-rate capacity against the array size for both the CAPA and SPDA in Figure 6(a).
+It is evident that the CAPA outperforms the SPDA for arbitrary array size.
 As anticipated, increasing the aperture size causes the sum-rate capacity to approach its upper bound, as derived in (67) and (72).
@@
 This verifies our discussions in Remark 9.
+Moreover, as the aperture size of the CAPA grows, the sum-rate achieved by the ZF detector asymptotically approaches the sum-rate capacity, which is consistent with the results in (68).
+This implies that the linear ZF detector, despite its lower implementation complexity, can deliver performance comparable to the optimal SIC scheme in the large-scale CAPA regime.
+To compare the performance gap between CAPA and the conventional SPDA in terms of sum-rate capacity, we plot the sum-rate capacity for SPDAs as a function of the array occupation ratio $\zeta_{\mathsf{oc}}$ in Figure 6(b).
+As shown in the graph, the sum-rate capacity for an SPDA gradually converges with that of a CAPA as the array occupation ratio increases towards $1$.
+This observation validates the statements in Remark 10.
 (a) Capacity region vs. $A$
-(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$.
+(b) Capacity regions with MC, $A=1$m2 and $d=\frac{\lambda}{3}$.
 Figure 7: Uplink capacity regions.
-Figure 6 shows the uplink capacity region achieved by the CAPA, where the green and blue points represent the achieved rates by SIC decoding order $1\rightarrow 2$ and $2\rightarrow 1$, respectively.
+Figure 7 shows the uplink capacity region achieved by the CAPA, where the two corner points represent the achieved rates by SIC decoding order $1\rightarrow 2$ and $2\rightarrow 1$.
 The rate tuple on the line segment connecting these two points is achieved by the time-sharing strategy.
@@
 This phenomenon occurs because the IUI decreases as the aperture size increases, which is consistent with the results shown in Figure 5.
-Furthermore, as depicted in Figure 6(b), the capacity region of the CAPA surpasses that of the SPDA, and their disparity narrows as the array occupation rate $\zeta_{\mathsf{oc}}$ increases, reinforcing findings from Figure 5(a).
-These findings underscore the superiority of CAPAs over conventional SPDAs in terms of channel capacity.
-Figure 7: Downlink transmission rates achieved by CAPA.
+Figure 8: Downlink transmission rates of CAPA.
 (a) Sum-rate capacity vs. $A$.
@@
 Figure 9: Downlink sum-rate capacity.
+To further examine the impact of MC, Figure 7(b) illustrates the rate regions with and without MC for both the CAPA and the SPDA.
+Since an exact MC model for continuous apertures remains an open research question, we approximate the CAPA as an SPDA with densely packed, edge-to-edge antenna elements [48], where the element size is set as $\sqrt{A_{\mathsf{s}}}=d=\frac{\lambda}{3}$.
+In this case, the coupling matrix is modeled as $\left(z_{a}+z_{t}\right)\left(\mathbf{Z}+z_{t}\mathbf{I}\right)^{-1}$, where $z_{a}$ represents the antenna impedance, $z_{t}$ denotes the termination impedance of each element, and $\mathbf{Z}$ is the mutual impedance matrix [39].
+We set $z_{a}=z_{t}=50~{}\Omega$ and model the mutual impedance elements as $\left[\mathbf{Z}\right]_{i,j}=0.1\frac{\mathrm{e}^{-\mathrm{j}k_{0}d_{ij}}}{d_{ij}^{2}}$, where $d_{ij}$ represents the distance between elements.
+From Figure 7(b), we observe that MC significantly degrades the performance of both CAPA and SPDA systems.
+However, despite the performance degradation caused by MC, the capacity region of CAPA remains larger than that of SPDA, highlighting its inherent advantage in terms of channel capacity.
+(a) CAPA capacity region vs. $A$.
+(b) Capacity regions with MC, $A=1$m2 and $d=\frac{\lambda}{3}$.
+Figure 10: Downlink capacity regions.
 VI-B Downlink CAPA Communications
-Figure 7 plots the per-user rate and the sum-rate capacity in terms of the total transmit SNR $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}$ in downlink CAPA communications.
-It is observed that the transmission rates of both users exhibit identical high-SNR slopes, regardless of the DPC encoding orders.
+Figure 8 plots the per-user rates, the sum-rate capacity, and the sum-rate achieved by the ZF precoder as functions of the aperture size in downlink CAPA communications.
+A notable performance gap is observed between the sum-rate achieved by the linear ZF precoder and the sum-rate capacity.
 Figure 9(a) and Figure 9(b) illustrate the downlink sum-rate capacity with respect to the aperture size and array occupation ratio, respectively.
-Similar to the uplink case, the downlink sum-rate capacity converges to its upper bound as the aperture size increases.
-Additionally, for the SPDA, the achieved sum-rate capacity is positively correlated with the array occupation ratio, further suggesting that the CAPA outperforms the SPDA in terms of channel capacity.
-(a) CAPA capacity region vs. $A$.
-(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$
-Figure 9: Downlink capacity regions.
+Similar to the uplink case, as the aperture size increases, the downlink sum-rate capacity converges to its upper bound, and the sum-rate achieved by the ZF scheme converges to the capacity.
+Additionally, for the SPDA, the achieved sum-rate capacity is positively correlated with the array occupation ratio, further suggesting that the CAPA outperforms the SPDA in terms of channel capacity..
 Figure 10 illustrates the downlink capacity region achieved by the CAPA.
@@
 This means that the boundary of the downlink capacity region is formed by the corner points of the dual uplink capacity regions, as illustrated in Figure 10(a).
-It is observed that the downlink capacity region of the CAPA expands with the aperture size $A$, whereas the capacity region of the SPDA expands with the array occupation ratio $\zeta_{\mathsf{oc}}$.
-Taken together, these observations demonstrate that the CAPA achieves a more extensive downlink capacity region than the conventional SPDA, highlighting its superiority.
+It is also observed that the downlink capacity region of the CAPA expands with the aperture size $A$.
+In Fig.
+10(b), the MC effects are demonstrated using the same method as in the uplink case.
+Additionally, in downlink scenario, the CAPA also achieves a more extensive capacity region than the conventional SPDA, highlighting its superiority.
 VII Conclusion
-This article has proposed an analytically tractable framework for CAPA-based wireless communications by utilizing continuous operators to address the continuous nature of the EM field.
+This article has proposed an analytically tractable framework for CAPA-based multiuser communications by utilizing continuous operators to address the continuous nature of the EM field.
 This framework significantly differs from the conventional matrix-based transmission framework for SPDA-based communications.
 Based on this novel framework, we have characterized the capacity limits of uplink and downlink channels for CAPAs in both single-user and multiuser cases.
-For each investigated scenario, we derived closed-form expressions for the capacity-achieving detectors or source currents along with the corresponding capacity.
-Additionally, we have analyzed the channel capacity by specializing the derived results to several typical array structures.
+For each scenario, we derived closed-form expressions for the capacity-achieving detectors or source currents along with the corresponding capacity.
+Additionally, we analyzed the channel capacity by specializing the derived results to several typical array structures.
 Through both theoretical analyses and numerical simulations, we have demonstrated that CAPA exhibits a higher sum-rate capacity and a more extensive capacity region compared to the conventional SPDA.
 This underscores CAPA’s potential as a promising transmission paradigm for future wireless networks.
+Acknowledgement
+The authors would like to thank the editor, Dr. Amine Mezghani, and anonymous reviewers for helpful comments that greatly improved the quality of the article.
 -A Proof of Lemma 1
@@
 Recalling that $\mu_{2}=-\frac{1}{{\mathsf{g}}_{2}}\pm\frac{1}{{\mathsf{g}}_{2}\sqrt{1+\overline{\gamma}_{\mathsf{dl},2}{\mathsf{g}}_{2}}}$ yields $\frac{2\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}}-(\frac{\mu_{2}}{1+\mu_{2}{\mathsf{g}}_{2}})^{2}{\mathsf{g}}_{2}=-\overline{\gamma}_{\mathsf{dl},2}$, which yields
-$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}.$
+$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}.$ (A42)
 As proved above, we have ${\mathsf{P}}_{1}\leq\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}$, and thus the powers satisfy
@@
 $\mathsf{g}_{k}=\int_{\frac{L_{z}}{-2}}^{\frac{L_{z}}{2}}{\int_{\frac{L_{x}}{-2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{k}^{*}(x,z)\mathsf{Q}_{k}(x,z)\mathrm{d}x}\mathrm{d}z}=\frac{r_{k}\Psi_{k}}{4\pi}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}$
-×$\times{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}}{{(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{-\frac{3}{2}}\mathrm{d}x}\mathrm{d}z}.$
-Subsequently, we can calculate the inner integral with the aid of [25, Eq.
-(2.264.5)] and the outer integral with the aid of [25, Eq.
+×$\times{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}}{{(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{-\frac{3}{2}}\mathrm{d}x}\mathrm{d}z}.$ (A45)
+Subsequently, we can calculate the inner integral with the aid of [56, Eq.
+(2.264.5)] and the outer integral with the aid of [56, Eq.
 (2.284)], which yields the results in (64).
 The channel correlation factor can be written as follows:
-$\rho=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{1}^{*}(x,z)\mathsf{Q}_{2}(x,z)\mathrm{d}x}\mathrm{d}z}.$ (A44)
-The above integrals can be calculated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\sum_{j=1}^{n}{f\left(x_{j}\right)}$ with $x_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$, lead to the results in (61).
+$\rho=\frac{1}{\sqrt{\mathsf{g}_{1}\mathsf{g}_{2}}}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{1}^{*}(x,z)\mathsf{Q}_{2}(x,z)\mathrm{d}x}\mathrm{d}z}.$ (A46)
+The above integrals can be calculated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\frac{\pi}{n}\sum_{j=1}^{n}{f\left(x_{j}\right)}$ with $x_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$, leading to the results in (65).
 This completes the proof of Lemma 4.
@@
 It follows from the concept of double integral that
-$\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}\!\!\!{f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)\epsilon_{k}^{2}}\approx\!\!\iint_{{\mathcal{C}}_{k}}{f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$
+$\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}\!\!\!{f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)\epsilon_{k}^{2}}\approx\!\!\iint_{{\mathcal{C}}_{k}}{f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$ (A49)
 Consequently, (A48) can be rewritten as follows:
 $\begin{split}\mathsf{g}_{k}\approx\frac{\zeta_{\mathsf{oc}}\Psi_{k}}{4\pi}\int_{-\frac{M_{z}\epsilon_{k}}{2}}^{\frac{M_{z}\epsilon_{k}}{2}}{\int_{-\frac{M_{x}\epsilon_{k}}{2}}^{\frac{M_{x}\epsilon_{k}}{2}}}f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z,\end{split}$ (A50)
-which can be calculated by utilizing [25, Eqs. (2.264.5) & (2.284)]. Furthermore, the expression of $\rho_{\mathsf{s}}$ follows from its definition, which completes the entire proof.
+which can be calculated by utilizing [56, Eqs. (2.264.5) & (2.284)]. Furthermore, the expression of $\rho_{\mathsf{s}}$ follows from its definition, which completes the entire proof.